### v2

# Weaving Geodesics and New Phenomena in Horocyclic Dynamics

Abstract.
We construct geometrically infinite hyperbolic surfaces supporting horocycles with tailored recurrence properties. In particular, we obtain the first examples of non-trivial minimal horocyclic orbit closures and of infinite locally-finite conservative horocyclic invariant measures which are singular with respect to the geodesic flow. Other examples include surfaces supporting horocyclic orbit closures of arbitrary Hausdorff dimension in $(1,2)$.
The study of horospherical flows on hyperbolic manifolds dates back to Hedlund in the 1930s [Hed36] and has played an important role in the development of modern homogeneous dynamics. In the finite-volume (and geometrically finite) cases, both the measure-theoretic and topological properties of the flow have been extensively studied, revealing a remarkable degree of rigidity (see, e.g. [Fur73, DS84, Bur90, Rob03, Rat91]). In contrast, the behavior of horospherical flows in the general geometrically infinite setting remains much less well understood.
Until recently the only explicitly described examples of horocyclic orbit closures on orientable hyperbolic surfaces were “trivial” — either the full non-wandering set for the horocyclic flow or single closed horocycles. While the existence of other, more intricate, orbit closures was known for decades, none have been described in detail, leaving much mystery as to their potential regularity and rigidity properties (c.f. [DS00, CM10, GL17, Mat16, Bel18b, Led97, Led98]).
In recent works [FLM23, FLM24], an explicit description of all horocyclic orbit closures was given in the setting of $\mathbb{Z}$-covers of compact hyperbolic surfaces. These orbit closures were shown to be highly irregular; their structural features depend in a delicate way on the geometry of the underlying surface. Intriguingly, all had integer Hausdorff dimension. They are also non-minimal.
Notably, no non-trivial minimal subsets for the horocyclic flow were known before now.
In this paper, we provide the first examples of non-trivial minimal horocyclic subsets as well as new fractional dimensional orbit closures.
As a consequence of our constructions, we provide the first counterexamples to the horospherical infinite measure rigidity phenomenon, which has been observed in a vast variety of settings, where every horospherically invariant ergodic Radon measure is either quasi-invariant under the geodesic flow (à la Babillot–Ledrappier) or supported on a single closed orbit; see [Bur90, Rob03, Sar04, LS07, Sar10, OP17, Sar19, LL22, Lan21, LLLO23] and Remark 2.8.
Main Results
Let $\Sigma$ be any orientable hyperbolic surface with unit tangent bundle $\mathsf{T}^{1}\Sigma\cong G/\Gamma$, where $G=\operatorname{PSL}_{2}(\mathbb{R})$ and $\Gamma\leq G$ is a discrete torsion-free subgroup acting isometrically on the right. Let $A=\{a_{t}=\mathrm{diag}(e^{t/2},e^{-t/2})\}_{t\in\mathbb{R}}$ denote the diagonal subgroup of $G$ generating, via left multiplication, the geodesic flow. Let $A_{+}=\{a_{t}:t\geq 0\}$, and let $N\leq G$ be the lower unipotent subgroup corresponding to the stable horocyclic flow on $\mathsf{T}^{1}\Sigma$. We denote by $p$ the projection map from the unit-tangent bundle (of either $\mathbb{H}^{2}$ or $\Sigma$) down to the surface.
Given a discrete subgroup $\Gamma\leq G$, we denote by $\Lambda\subseteq\partial\mathbb{H}^{2}$ its limit set. The non-wandering set for the horocycle flow is
$\mathcal{E}=\{g\Gamma\in G/\Gamma:g^{+}\in\Lambda\},$
where $g^{+}$ is the terminal endpoint in $\partial\mathbb{H}^{2}$ of the geodesic ray emanating from $g$.
Recall that a non-empty $N$-invariant closed set $F\subseteq\Sigma$ is called $N$-minimal if all $N$-orbits in $F$ are dense in $F$. A characterization of points with dense horocyclic orbits in $\mathcal{E}$ is given by [Ebe77, Dal00] where it was shown that $\overline{Nx}\neq\mathcal{E}$ if and only if the geodesic ray $A_{+}x$ is quasi-minimizing, that is, $d_{\mathsf{T}^{1}\Sigma}(a_{t}x,x)\geq t-c$ for some $c\geq 0$ and every $t\geq 0$. As a consequence, $\mathcal{E}$ is $N$-minimal if and only if $\Gamma$ is convex co-compact.
Studying the different possible trajectories of quasi-minimizing rays and their “efficiency” has turned out to be key in the analysis of horocyclic orbit closures. Drawing on techniques developed in [FLM23, FLM24] and inspired by examples introduced by Alexandre Bellis in [Bel18a, §1.5.1], we provide a recipe for tailoring geometrically infinite surfaces supporting horocycles with prescribed recurrence properties. Our main results are the following:
Theorem.
  1. (1)
    There exists a surface $\Sigma$ such that $\mathsf{T}^{1}\Sigma$ supports an $N$-minimal closed subset which is neither $\mathcal{E}$ nor a single $N$-orbit. Moreover, this minimal orbit closure supports an $N$-invariant, ergodic, infinite and locally finite measure $\mu$ which is conservative but singular with respect to the geodesic flow, that is, $a_{t}.\mu\perp\mu$ for all $t\neq 0$.
  2. (2)
    For any $\alpha\in(1,2)$ there exists a surface $\Sigma_{\alpha}$ such that $\mathsf{T}^{1}\Sigma_{\alpha}$ supports an $\alpha$-Hausdorff dimensional horocyclic orbit closure.
Remarks.
  • Our surfaces are extremely sparse; the injectivity radius along all diverging geodesic rays tends to infinity. This implies, in particular, that the tameness conditions imposed in [Sar10] to deduce measure rigidity cannot be removed. Equivalent geometric conditions appear in [LL22].
  • The possible non-regularity of orbit closures we construct is quite extreme, allowing us to construct orbit closures having disagreeing Hausdorff and lower/upper Minkowski dimensions; see §3.1.2.
  • Note that for any $\lambda\in(0,1]$ there exist convex co-compact Fuchsian groups having $\lambda$-dimensional limit sets. In such surfaces, the corresponding orbit closure $\mathcal{E}$, being $AN$-invariant, is hence $2+\lambda$-dimensional. We may thus conclude that any $\alpha\in[1,3]$ can be the dimension of some horocycle orbit closure.
Remark to the reader about the proof
While we rely on techniques developed in [FLM23, FLM24], we will only make use of several elementary insights and lemmas from said papers. Our proof is fairly self-contained and requires no prior knowledge or understanding of the results in the $\mathbb{Z}$-cover setting.
1. Setup
1.1. Loom Surfaces
It will be convenient for us to work with the band model for the hyperbolic plane, that is, the space $\mathbb{H}^{2}:=\{z\in\mathbb{C}:|\mathrm{Im}z|<\pi/2\}$ equipped with the metric $|dz|/\cos\mathrm{Im}z$.
Given a closed convex domain $J\subset\mathbb{H}^{2}$ with totally geodesic boundary we denote by $\hat{J}$ its double, that is, the space
$\widehat{J}=\overline{J}\times\{0,1\}/\sim\qquad\text{where}\quad(z,0)\sim(z,1)\quad\text{for all }z\in\partial J.$
Under these conditions, $\widehat{J}$ is a complete hyperbolic surface without boundary.
For $s\in\mathbb{R}$ and $h\in(0,\pi/2)$ we denote by $D_{h}(s)$ the unique open half-plane contained in $\{\mathrm{Im}z>0\}\cap\mathbb{H}^{2}$ and bounded by the geodesic which is perpendicular to $s+(-\pi/2,\pi/2)i$ at the point $s+hi$, see Figure 1.
Figure 1. Different geodesics in the band model (in blue and red). A half plane $D_{h}(s)$ (shaded in gray).
Given a sequence $s=(s_{j},h_{j})_{j\in\mathbb{N}}$ satisfying $\overline{D_{h_{k}}(s_{k})}\cap\overline{D_{h_{j}}(s_{j})}=\emptyset$ for all $k\neq j$, we consider the surface
$\Sigma_{s}=\widehat{J_{s}}\quad\text{where}\quad J_{s}=\mathbb{H}^{2}\smallsetminus\bigcup_{k}D_{h_{k}}(s_{k}).$
Let $q:\Sigma_{s}\to\overline{J_{s}}\subset\mathbb{H}^{2}$ be the quotient mapping identifying the two copies of $J_{s}$ comprising $\Sigma_{s}$. Topologically, $\Sigma_{s}$ is a plane with a countable discrete set of punctures, see Figure 2.
Definition 1.1.
A loom surface is a surface $\Sigma_{s}$ as above, where the sequence $s=(s_{k},h_{k})_{k\in\mathbb{N}}$ has $s_{k}$ monotonic increasing, $h_{k}$ bounded above by $c<\pi/2$, and satisfying
$d_{\mathbb{H}^{2}}(\partial D_{h_{k}}(s_{k}),\partial D_{h_{k+1}}(s_{k+1}))\to\infty\quad\text{where}\quad k\to\infty.$
Note that under the above conditions the function $\text{Inj--rad}:\Sigma_{s}\to(0,\infty)$, assigning the injectivity radius at a point, is a proper map. In particular, injectivity radius tends to infinity along any diverging geodesic ray.
Figure 2.
1.2. Tight map and stretch lamination
Consider $\Sigma_{s}$ as above and denote by $x_{j}$, $j=0,1$, the points in $\mathsf{T}^{1}\Sigma_{s}$ corresponding to $(0,j)\in J_{s}\times\{j\}$ with horizontal unit vector $1\in S^{1}$. Let $\ell\subset J_{s}\subset\mathbb{H}^{2}$ be the image of $p(Ax_{0})$ or $p(Ax_{1})$ under the quotient mapping $q$. We identify $\mathbb{R}\cong\ell$ by the rule $t\mapsto q(p(a_{t}x_{j}))$, and define $\tau:\Sigma_{s}\to\mathbb{R}$ as the composition of $q$ followed by the nearest point projection to $\ell$. [^1] The nearest point projection to $\ell$ in $J_{s}$ is the restriction of the nearest point projection to $\ell$ in $\mathbb{H}^{2}$, as $\ell\subset J_{s}$ and $J_{s}$ is geodesically convex. This map is $1$-Lipschitz and a strict contraction away from $\ell$.
As the composition of $1$-Lipschitz maps, $\tau$ is itself $1$-Lipschitz. Let $\lambda=p(Ax_{0})\cup p(Ax_{1})$ and $\mathsf{T}_{+}^{1}\lambda=Ax_{1}\cup Ax_{2}$. Observe that $\tau$ is strictly contracting away from $\lambda$ and is isometric along each of its components. In particular, $p(Ax_{0})$ and $p(Ax_{1})$ are isometrically embedded in $\Sigma$, and $Ax_{0}$ and $Ax_{1}$ are isometrically embedded in $\mathsf{T}^{1}\Sigma$.
Abusing notation, we also use $\tau$ to denote $p^{*}\tau:\mathsf{T}^{1}\Sigma_{s}\to\mathbb{R}$, which is constant along fibers of $p$. For $y\in\mathsf{T}^{1}\Sigma_{s}$, $\tau(y)=t$ means that the closest point to $p(y)$ in $\lambda$ is $p(a_{t}x_{0})$ or $p(a_{t}x_{1})$. Equivalently, $\tau(y)=\mathrm{Re}(q(p(y))$, where $\mathrm{Re}(z)$ is the real part of the complex number $z$.
1.3. Slack
We consider the notion of $\tau$-slack introduced in [FLM24, §3] which measures how “efficiently” a path progresses towards the end at $\infty$:
Definition 1.2.
Let $\alpha:[a,b]\to\Sigma_{s}$ be a rectifiable curve. We define the slack of $\alpha$ to be
$\mathscr{S}_{+}(\alpha)=\mathrm{length}(\alpha)-(\tau(\alpha(b))-\tau(\alpha(a))).$
Similarly if $\beta:[a,b]\to\mathsf{T}^{1}\Sigma_{s}$ is rectifiable we define its slack to be the slack of its projection to $\Sigma_{s}$. Note that $\mathscr{S}_{+}$ is non-negative and additive under concatenation of paths, so if $I\subset\mathbb{R}$ is connected and $\alpha:I\to\mathsf{T}^{1}\Sigma_{s}$, we can define
$\mathscr{S}_{+}(\alpha)=\lim_{T\to\infty}\mathscr{S}_{+}\left(\alpha|_{I\cap[-T,T]}\right)=\sup_{T>0}\mathscr{S}_{+}\left(\alpha|_{I\cap[-T,T]}\right)\in[0,\infty].$
If $\alpha$ is a geodesic flow line, of the form $A_{[s,t]}z$, we note that $\mathscr{S}_{+}(\alpha)$ is just $(t-s)-(\tau(a_{t}z)-\tau(a_{s}z))$ and that
(1.1) $\mathscr{S}_{+}(\alpha)=0\text{ if and only if }\alpha\subset\mathsf{T}^{1}_{+}\lambda.$
1.4. Busemann-type Function
Consider the function $\beta:\mathsf{T}^{1}\Sigma_{s}\to[-\infty,\infty)$ defined by
$\beta(y)=\tau(y)-\mathscr{S}_{+}(A_{+}y)=\lim_{t\to+\infty}\tau(a_{t}y)-t.$
It is upper semi-continuous, as a decreasing limit of continuous functions, and also $N$-invariant, see [FLM23, Lemma 6.1]. Therefore, for all $y\in\mathsf{T}^{1}\Sigma_{s}$ we have
(1.2) $\overline{Ny}\subseteq\beta^{-1}([\beta(y),\infty)).$
In particular, since $\tau(x_{j})=0$ and $\mathscr{S}_{+}(A_{+}x_{j})=0$ for both $j=0,1$, we conclude:
Fact 1.3.
For both $j=0,1$, all $y\in\overline{Nx_{j}}$ satisfy $\beta(y)\geq 0$.
1.5. Weaving Lemma
A fundamental observation is the following: any geodesic trajectory $A_{+}y$ spending an infinite amount of time a definite distance away from the isometric locus of $\tau$, $\mathsf{T}^{1}_{+}\lambda$, will necessarily have $\beta(y)=$-$\infty$. In other words,
Fact 1.4.
All $y\in\mathsf{T}^{1}\Sigma_{s}$ with $\beta(y)>$-$\infty$ satisfy that for all $\varepsilon>0$ there exists $T>0$ such that $d(a_{t}y,\mathsf{T}^{1}_{+}\lambda)<\varepsilon$ for all $t>T$.
See the first part of the proof of [FLM23, Thm. 3.4] for more details.
In the case of loom surfaces, this asymptotic behavior ensures all finite slack geodesic rays eventually follow some weaving pattern.
Given $k\in\mathbb{N}$ we denote by $\eta_{k}^{+}$ the unique geodesic in $\Sigma_{s}$ which is backward asymptotic to $A_{-}x_{0}$ and forward asymptotic to $A_{+}x_{1}$ crossing once from $J_{s}\times\{0\}$ to $J_{s}\times\{1\}$ through $\partial D_{h_{k}}(s_{k})$. We call $\eta_{k}^{+}$ a crossing. We similarly have $\eta_{k}^{-}$, the symmetric geodesic passing from $J_{s}\times\{1\}$ to $J_{s}\times\{0\}$ through $\partial D_{h_{k}}(s_{k})$, see Figure 3.
Figure 3.
Notice that by the symmetry of the construction we have $\eta_{k}^{\pm}\cap\partial D_{h_{k}}(s_{k})=\{s_{k}+ih_{k}\}$.
Definition 1.5.
A weaving pattern is a any subset of $\mathbb{N}$ thought of as a strictly increasing sequence of indices $W=\{k_{1}<k_{2}<$...\}.
A weaving geodesic ray with weaving pattern $W$ is an infinite geodesic ray $A_{+}w$ given by pulling tight the following concatenation:
$\alpha_{0}\ast\eta^{\sigma_{1}}_{k_{1}}\ast\eta^{\sigma_{2}}_{k_{2}}\ast\eta^{\sigma_{3}}_{k_{3}}$...
where:
  1. (1)
    $\sigma_{k}\in\{+,-\}$ are alternating signs;
  2. (2)
    $\alpha_{0}$ is the shortest segment connecting $p(w)$ to $\eta^{\sigma_{1}}_{k_{1}}$;
  3. (3)
    the $\ast$ notation is interpreted as concatenation of subsegments beginning and ending at the unique intersection points of the paths;
  4. (4)
    the weaving pattern is allowed to be finite in which case we append, at the end, an infinite subray of either $A_{+}x_{0}$ or $A_{+}x_{1}$, as needed.
Given a path $\alpha$, we denote by $\hat{\alpha}$ the geodesic segment given by pulling $\alpha$ tight, that is, the unique geodesic segment in $\alpha$’s homotopy class relative its endpoints (or endpoints at infinity). Given a weaving pattern $W$ we define
$\eta_{W,+}=\widehat{\eta_{k_{1}}^{+}\ast\eta_{k_{2}}^{-}\ast\eta_{k_{3}}^{+}\ast...}$,
where the alternating signs begin with a ++. We similarly define $\eta_{W,-}$. We will refer to such $\eta_{W}^{\pm}$ as weaving geodesics with weaving pattern $W$.
Figure 4. Intersection point of two crossings.
Lemma 1.6 (Weaving Lemma).
In a loom surface $\Sigma_{s}$, for any $\rho\geq 0$ there exists $S>0$ such that any geodesic ray $A_{+}y$ with slack $\mathscr{S}_{+}(A_{+}y)\leq\rho$ and beginning at $\tau(y)>S$ is weaving.
Proof.
First note that since both $J_{s}\times\{j\}$ are convex and simply connected, any geodesic ray passing from $J_{s}\times\{j\}$ through $\partial D_{h_{k}}(s_{k})$ to $J_{j+1}$ will either stay indefinitely in the $j+1\mod 2$-side or will have its first return to $J_{s}\times\{j\}$ happen through a different component of $\partial J_{s}$. Therefore, by the uniqueness of geodesic representatives in each homotopy class, it’s enough to show that the sequence of boundary components of $J_{s}$ which the geodesic crosses has increasing indices.
Now for any $\rho\geq 0$ there exists $k_{0}\in\mathbb{N}$ such
$\rho<\min_{k\geq k_{0}}d_{\mathbb{H}^{2}}(\partial D_{h_{k}}(s_{k}),\partial D_{h_{k-1}}(s_{k-1})).$
[…]
A direct consequence of 1.6 is that any finite slack geodesic ray is eventually weaving.
1.6. Slack of a Weaving Geodesic
As a first step we note that the slack of a single crossing is entirely determined by the “height” of the boundary component $h_{k}$. In fact using the hyperbolic cosine law one can verify the following:
Fact 1.7.
$\mathscr{S}_{+}(\eta_{k}^{\pm})=2\ln\cosh(d_{\mathbb{H}^{2}}(0,ih_{k}))=2\ln\cosh\left(\int_{0}^{h_{k}}\frac{dt}{\cos(t)}\right)$.
We invoke the following lemma from [FLM24, Lemma 3.5]:
Lemma 1.8.
For all $c>0$ there exist constants $\kappa_{c},\varepsilon_{0}>0$ such that the following holds for all $0<\varepsilon<\varepsilon_{0}$. Let $\alpha_{i}:[a_{i},b_{i}]\to\Sigma$ for i=1,,ni=1,...,n be a sequence of geodesic arcs, each of length greater or equal to $c$, and satisfying
(1.3) $\sum_{i=1}^{n-1}d_{\mathsf{T}^{1}\Sigma}(\mathsf{T}^{1}\alpha_{i}(b_{i}),\mathsf{T}^{1}\alpha_{i+1}(a_{i+1}))<\varepsilon$
and let $\bar{\alpha}$ denote an arc obtained from $\cup\alpha_{i}$ by joining each endpoint $\alpha_{i}(b_{i})$ to $\alpha_{i+1}(a_{i+1})$ using arcs whose total length is less than $\varepsilon$. Then there exists a geodesic arc $\alpha$ homotopic, relative to the endpoints, to $\bar{\alpha}$ and satisfying
$\left|\mathscr{S}_{+}(\alpha)-\sum_{i=1}^{n}\mathscr{S}_{+}(\alpha_{i})\right|<\kappa_{c}\cdot\varepsilon.$
Moreover, the Hausdorff distance between $\alpha$ and $\bar{\alpha}$ is smaller than $\kappa_{c}\varepsilon$.
The claim further holds with $\alpha_{n}:[a_{n},\infty)\to\Sigma$ and where $\alpha$ is a geodesic ray from $\alpha_{1}(a_{1})$ which is forward-asymptotic to $\alpha_{n}$; and similarly with $\alpha_{1}:(-\infty,b_{1}]\to\Sigma$.
The fact that the $h_{k}$’s are uniformly bounded above, together with the assumption that $d_{\mathbb{H}^{2}}(\partial D_{h_{k}}(s_{k}),\partial D_{h_{k+1}}(s_{k+1}))\to\infty$ implies $|s_{k}-s_{l}|\to\infty$ if $k\not$$\ell$ and $k,\ell\to\infty$ and that corresponding crossings $\eta_{k}^{\pm}$ and $\eta_{l}^{\mp}$ intersect at angles tending to 0. Moreover, taking increasing and exhausting portions of the crossing will account for all of $\mathscr{S}_{+}(\eta_{k}^{\pm})$, e.g. $\mathscr{S}_{+}(\eta_{k}^{\pm}|_{[-T,T]})\to\mathscr{S}_{+}(\eta_{k}^{\pm})$. We thus conclude the following:
Proposition 1.9.
For any $\varepsilon>0$ and $m\in\mathbb{N}$, there exists $S>0$ such that any length $m$ weaving pattern, $W=\{k_{1}<k_{2}<...<k_{m}\}$, satisfying
$|s_{k_{j}}-s_{k_{j+1}}|>S\quad\text{for all}\quad 1\leq j<m$
will have:
$\left|\mathscr{S}_{+}(\eta_{W,\pm})-\sum_{j=1}^{m}\mathscr{S}_{+}\left(\eta_{k_{j}}^{\pm}\right)\right|<\varepsilon.$
A fundamental observation, highlighting the utility of the notion of slack, is the following lemma proven in [FLM24, Lemma 3.3]:
Lemma 1.10.
For any $x,y\in\mathsf{T}^{1}_{+}\lambda$ with $\tau(x)=\tau(y)$ and $t\geq 0$, we have $a_{t}y\in\overline{Nx}$ if and only if there exists $y_{m}\to y$ such that $A_{+}y_{m}$ is asymptotic to $A_{+}x$, and $\mathscr{S}_{+}(A_{+}y_{m})\to t$.
The elementary proof of this lemma applies verbatim to the setting of this paper. We will make explicit use of this lemma in Section 3.
2. Minimal Orbit Closures
In this section we will work under the following assumption:
Definition 2.1.
A loom surface $\Sigma_{s}$ with $s=(s_{j},h_{j})_{j\in\mathbb{N}}$ is said to satisfy the summability condition if
$\sum_{j=1}^{\infty}\mathscr{S}_{+}(\eta_{j}^{\pm})=\sum_{j=1}^{\infty}2\ln\cosh(d_{\mathbb{H}^{2}}(0,ih_{k}))<\infty.$
We prove the following:
Theorem 2.2.
Let $\Sigma_{s}$ be a loom surface satisfying the summability condition, then the orbit closure $F=\overline{Nx_{0}}$ is $N$-minimal and satisfies
$Nx_{0}\subsetneq F\subsetneq\mathcal{E},$
where $\mathcal{E}$ denotes the non-wandering set for the horocycle flow on $\mathsf{T}^{1}\Sigma_{s}$.
Recall that two points $y,w\in\mathsf{T}^{1}\Sigma_{s}$ are called $A$-proximal if
$\liminf_{t\to+\infty}d(a_{t}y,a_{t}w)=0.$
We will make use of the following fact, proven in [FLM23, Cor. 8.3]:
Proposition 2.3.
If $y,w\in\mathsf{T}^{1}\Sigma_{s}$ are $A$-proximal then $\overline{Ny}=\overline{Nw}$.
Our strategy for proving Theorem 2.2 will be to show that all the points in $\overline{Nx_{0}}$ are $A$-proximal to $x_{0}$. Two key observations under these assumptions:
Lemma 2.4.
If $\beta(y)=b>$-$\infty$ then $y$ is $A$-proximal to $a_{b}x_{0}$.
Proof.
First, note that the summability condition implies in particular that $h_{k}\to 0$ and hence
(2.1) $\lim_{k\to\infty}d\left(a_{s_{k}}x_{0},a_{s_{k}}x_{1}\right)=0.$
In particular, for all $k$ large enough we have
(2.2) $\mathsf{T}^{1}_{+}\lambda\cap\tau^{-1}\left((s_{k}-\varepsilon,s_{k}+\varepsilon)\right)\subset B_{3\varepsilon}(a_{s_{k}}x_{0}).$
Assume $y$ as in the statement. For any $\varepsilon>0$ there exists $T>0$ large enough such that $a_{t}y$ is in the $\varepsilon$-neighborhood of $\mathsf{T}^{1}_{+}\lambda$ for all $t\geq T$ (1.4). By the definition of $\beta$ we have $|\tau(a_{t-b}y)-t|\to 0$ as $t\to\infty$. Hence for all $k$ large enough we have that $|\tau(a_{s_{k}-b}y)-\tau(a_{s_{k}}x_{0})|<\varepsilon$ and $a_{s_{k}-b}y$ is in the $\varepsilon$-neighborhood of $\mathsf{T}^{1}_{+}\lambda$. By (2.2) we conclude
$\liminf_{k\to\infty}d(a_{s_{k}-b}y,a_{s_{k}-b}a_{b}x_{0})\leq 3\varepsilon,$
implying the claim. ∎
The second observation is:
Lemma 2.5.
For any $\varepsilon>0$ there exists $k_{0}\in\mathbb{N}$ such that any weaving pattern $W$ satisfying $W\geq k_{0}$ will have $\mathscr{S}_{+}(\eta_{W,\pm})<\varepsilon$.
Proof.
By the summability condition, for any $\varepsilon>0$ there exists $k_{0}$ large enough for which
$\sum_{j=k_{0}}^{\infty}\mathscr{S}_{+}(\eta_{j}^{\pm})<\varepsilon.$
Recall that the past of $\eta_{W,\pm}$ is asymptotic to $\mathsf{T}^{1}_{+}\lambda$ and that this geodesic has arbitrarily long segments which are arbitrarily close to $\mathsf{T}^{1}_{+}\lambda$ (close to intersection points of the crossings). This is also true for the “untightened” path $\alpha_{W}=\eta_{k_{1}}^{\pm}\ast\eta_{k_{2}}^{\mp}\ast$.... Hence we can compare slacks along exhausting subsegments of $\eta_{W,\pm}$ and $\alpha_{W}$ which have $\tau$-value arbitrarily close at their respective endpoints. The main difference is that the path along $\eta_{W,\pm}$ is shorter. Therefore we have $\mathscr{S}_{+}(\eta_{W,\pm})<\mathscr{S}_{+}(\alpha_{W})$. The inequality
$\mathscr{S}_{+}(\alpha_{W})<\sum_{j=k_{0}}^{\infty}\mathscr{S}_{+}(\eta_{j}^{\pm})$
follows from the additivity and non-negativity of slack. ∎
Proof of Theorem 2.2.
First, by (2.1) we note that $x_{0}$ and $x_{1}$ are $A$-proximal. Therefore by 2.3 we have $x_{1}\in F=\overline{Nx_{0}}$. Since clearly $x_{1}\notin Nx_{0}$ (the corresponding rays $A_{+}x_{j}$ are not asymptotic) we deduce that $F\neq Nx_{0}$. The strict inclusion $F\subsetneq\mathcal{E}$ follows from $x_{0}$ being quasi-minimizing.
The main part of the claim is the $N$-minimality. Our strategy will be to show that all accumulation points of $Nx_{0}$ are $A$-proximal to $x_{0}$ therefore by 2.3 have dense $N$-orbits in $F$.
Let $y\in F$ be an accumulation point of $Nx_{0}$ and let $0<\varepsilon<1$ be arbitrary. By Facts 1.3 and 1.4 we know $\mathscr{S}_{+}(A_{+}y)<\infty$ and that for any $\varepsilon>0$ a subray of $A_{+}y$ is contained in the $\varepsilon$-neighborhood of $\mathsf{T}^{1}_{+}\lambda$. Let $T_{1}>0$ be such that $\mathscr{S}_{+}(A_{[T_{1},\infty)}y)<\varepsilon$ and $A_{[T_{1},\infty)}y$ is contained in the $\varepsilon$-neighborhood of $\mathsf{T}^{1}_{+}\lambda$.
[…]
Now let $nx_{0}$ be sufficiently close to $y$ so as to satisfy $d(a_{T_{2}}nx_{0},a_{T_{2}}y)<\varepsilon$, for $T_{2}$ above, and also $d(nx_{0},y)<1$. Set $w=a_{T_{2}}nx_{0}$. Recall that the function $\beta$ is $N$-invariant and hence $\beta(nx_{0})=\beta(x_{0})=0$ implying that $\mathscr{S}_{+}(A_{+}nx_{0})=\tau(nx_{0})$. Since $\tau$ is 1-Lipschitz we have $\tau(nx_{0})<\tau(y)+1$ and hence
$\mathscr{S}_{+}(A_{+}w)\leq\mathscr{S}_{+}(A_{+}nx_{0})<\tau(y)+1=\rho.$
By our choice of $T_{2}$ we conclude that $A_{+}w$ is weaving with pattern $W\geq k_{0}$. That is, $A_{+}w$ is asymptotic to some $\eta_{W,\pm}$ and moreover $d(w,\eta_{W,\pm})<2\varepsilon$. This implies that
(2.3) $\mathscr{S}_{+}(A_{+}w)<2\varepsilon+\mathscr{S}_{+}(\eta_{W,\pm})<3\varepsilon.$
But recall that $d(a_{t}nx_{0},a_{t}x_{0})\to 0$ as $t\to\infty$ and hence $d(a_{t-T_{2}}w,a_{t}x_{0})\to 0$. Joining this fact with (2.3) we deduce that $|\tau(w)-T_{2}|<3\varepsilon$.
Since $d(w,a_{T_{2}}y)<\varepsilon$ we have $|\tau(a_{T_{2}}y)-T_{2}|<4\varepsilon$. On the other hand, by our choice of $T_{2}>T_{1}$ we had
$\mathscr{S}_{+}(A_{[T_{2},\infty)}y)<\varepsilon.$
We thus conclude $|\beta(a_{T_{2}}y)-T_{2}|<5\varepsilon$, or in other words, $|\beta(y)|<5\varepsilon$ which by Lemma 2.4 implies $y$ is $A$-proximal to $a_{b}x_{0}$ for some $|b|<5\varepsilon$ (in fact, necessarily $0\leq b$). Since $\varepsilon$ was arbitrary we conclude the claim. ∎
Remark 2.6.
Note that minimality together with (1.2) implies that $\beta$ is constant (zero) on all of $\overline{Nx_{0}}$. This also follows directly from our proof. Moreover, the fact that $\beta(a_{t}z)=\beta(z)+t$ for all $t\in\mathbb{R}$ implies that $a_{t}y\notin\overline{Nx_{0}}$ for all $y\in\overline{Nx_{0}}$ and $t\neq 0$.
2.1. Invariant Measures
In this subsection we show the following:
Theorem 2.7.
Let $\Sigma_{s}$ be a loom surface satisfying the summability condition, then the minimal orbit closure $\overline{Nx_{0}}$ supports a locally finite $N$-invariant and ergodic measure, $\mu$, which is infinite, conservative and singular with respect to the geodesic flow, that is,
$a_{t}.\mu\perp\mu\quad\text{for all }t\neq 0.$
This theorem provides the first examples of $N$-ergodic and invariant locally finite measures which are neither quasi-invariant with respect to the geodesic flow nor supported on a single closed horocycle.
Proof of 2.7.
The existence of an $N$-invariant locally finite measure supported on $\overline{Nx_{0}}$ follows from [KMR13, Lemma 2.2] and the superamenability of the group $\mathbb{R}$ [Ros74]. See also [Ami22]. For completeness and accessibility, we provide a short, more hands-on, construction of such measures in A.1 in the appendix.
These statements ensure the existence of an $N$-invariant, conservative and locally-finite measure supported on $\overline{Nx_{0}}$. Let $\mu$ denote a locally-finite ergodic component of said measure. By Ratner’s classification of finite $N$-invariant measures [Rat92] (and the fact that $\overline{Nx_{0}}$ is not a homogeneous subspace of $\mathsf{T}^{1}\Sigma_{s}$) we conclude $\mu$ is necessarily infinite.
Singularity of $\mu$ with respect to the geodesic flow follows from the fact that $a_{t}\overline{Nx_{0}}\cap\overline{Nx_{0}}=\emptyset$ for all $t\neq 0$, see Remark 2.6, implying that the topological support of $\mu$ is not invariant under any $a_{t}$ with $t\neq 0$. ∎
Remark 2.8.
We note that the notion of a “trivial” measure in this paper is strictly broader than the one used by Sarig in [Sar10], who referred only to measures corresponding to periodic horocycles or orbits based outside the limit set. The measures constructed in [Sar10, Theorem 3] are trivial in the sense that they are supported on single proper horocycle orbits, which was known by the author.
The measures constructed in [Sar10, Theorem 3] are ergodic components of Lebesgue measure on certain surfaces given by Fuchsian groups of the first kind with critical exponent<1/2<\!1/2, see [Pat79]. In [Pat77, Equation (18)], Patterson shows that Fuchsian groups $\Gamma$ having critical exponent smaller than one-half satisfy, in the Poincaré disk model,
$\sum_{\gamma\in\Gamma}|\gamma^{\prime}(\xi)|<\infty\quad\text{for Lebesgue-almost every }\xi\in S^{1}.$
Recall that in the disk model one has $|\gamma^{\prime}(\xi)|=e^{-\beta_{\xi}(\gamma^{-1}.0,0)}$ where $\beta$ is the Busemann cocycle, see e.g. [Sar19, Appendix 1]. Therefore, we have that for Leb-a.e. $\xi\in\partial\mathbb{H}^{2}$, any horocycle tangent to $\xi$ has only finitely many points in $\Gamma.0$ within a bounded distance of it. This in turn implies that such horocycles have no accumulation points in the quotient surface.
3. Fractional Hausdorff Dimension
In this section we consider surfaces satisfying the following:
Definition 3.1.
A loom surface $\Sigma_{s}$ with $s=(s_{j},h_{j})_{j\in\mathbb{N}}$ is called distal if
$\inf_{j\in\mathbb{N}}h_{j}>0.$
Note that under these conditions, the two points $x_{0}$ and $x_{1}$ are $A$-distal, i.e. they satisfy $\inf_{t\to\infty}d(a_{t}x_{0},a_{t}x_{1})>0$.
We observe that 1.4 implies, in this case, that any finite slack geodesic ray is eventually asymptotic to either $A_{+}x_{0}$ or $A_{+}x_{1}$ since the $\varepsilon$-neighborhood of $\mathsf{T}^{1}_{+}\lambda$ is disconnected for small enough $\varepsilon$. In other words, all points in $\mathsf{T}^{1}\Sigma_{s}$ supporting finite slack rays are contained in one of two distinct $AN$-orbits $ANx_{0}\sqcup ANx_{1}$. The function $\beta$ is $N$-invariant and also, by the definition, $A$-equivariant in the sense that for all $y$ and $t\in\mathbb{R}$ we have $\beta(a_{t}y)=\beta(y)+t$. 1.3 hence implies:
Lemma 3.2.
If $\Sigma_{s}$ is a distal loom surface, then
$\overline{Nx_{0}}\subseteq A_{+}Nx_{0}\sqcup A_{+}Nx_{1}.$
In this section we are interested in the Hausdorff dimension of $\overline{Nx_{0}}$, we may thus consider each component of the orbit closure separately. Each $A_{+}N$-orbit in $\mathsf{T}^{1}\Sigma_{s}$ is a locally isometric projection of countably many $A_{+}N$-orbits in $\mathsf{T}^{1}\mathbb{H}^{2}$. Each lift of $\overline{Nx_{0}}\cap A_{+}Nx_{j}$ is a product set which is $N$-saturated, implying
$\dim_{\mathrm{H}}\overline{Nx_{0}}\cap A_{+}Nx_{j}=1+\dim_{\mathrm{H}}\overline{Nx_{0}}\cap A_{+}x_{j}.$
Hence our analysis boils down to understanding the sets
$\Delta_{j}=\{t\geq 0:a_{t}x_{j}\in\overline{Nx_{0}}\cap A_{+}x_{j}\}.$
See [FLM24, §2.2] and [FLM23, §7] for another account of such sets (under different notations).
The main theorem in this section is the following:
Theorem 3.3.
Let $\Sigma_{s}$ be a distal loom surface. Let $E=\mathrm{accum}\left(\mathscr{S}_{+}(\eta_{k}^{\pm})\right)_{k\in\mathbb{N}}$ be the accumulation points in $\mathbb{R}$ of the sequence of slacks of crossings. Then
(3.1) $\Delta_{0}=\bigcup_{k=1}^{\infty}2kE\qquad\text{and}\qquad\Delta_{1}=\bigcup_{k=0}^{\infty}(2k+1)E,$
where $mE:=\underbrace{E+\cdots+E}_{m}$.
Remark 3.4.
Note that the distality assumption ensures $E>\delta>0$ for some $\delta$ and hence for all $T>0$
$\bigcup_{m\in\mathbb{N}}mE\cap[0,T]\subseteq\bigcup_{m=1}^{\lceil\frac{T}{\delta}\rceil}mE,$
implying in particular that both $\Delta_{j}$ are closed.
The vast flexibility of our construction allows us to exhibit any compact subset of $(0,\infty)$ as $E$, that is, we have:
Proposition 3.5.
For any compact [^2] Our definition of loom surfaces is not the most general imaginable. In particular, the assumption that the $h_{k}$ are bounded is merely imposed for simplicity. One can construct surfaces with an unbounded sequence of $h_{k}$ under which additional conditions on $|s_{k+1}-s_{k}|\gg 0$ would ensure all the claims of this paper will hold. set $E\subset(0,\infty)$ there exists a distal loom surface $\Sigma_{s}$ in which $E=\mathrm{accum}\left(\mathscr{S}_{+}(\eta_{k}^{\pm})\right)_{k\in\mathbb{N}}$.
Proof.
Given $E$, choose a finite or countable dense subset $\{e_{1},e_{2},$...\} of $E$. By solving for $h$ in 1.7 we can construct $\Sigma_{s}$ where every $e_{j}$ appears infinitely many times as the slack of different crossings. The proposition follows. ∎
3.1. Examples
These results provide a rich source of interesting and non-regular examples. Here are a few:
3.1.1. Fixed Fractional Hausdorff Dimension
For any $\alpha\in[0,1]$ there exists a compact set $E^{\prime}$ with $\dim_{\mathrm{H}}mE^{\prime}=\alpha$ for all $m\geq 1$, see [SS10, Kör08]. Since $E^{\prime}$ may contain 0, and the accumulations of slacks of geodesics $\eta_{k}^{\pm}$ are all strictly positive in a distal loom surface, we take $E=1+E^{\prime}$. By Proposition 3.5, there is a distal loom surface with $E=\mathrm{accum}\left(\mathscr{S}_{+}(\eta_{k}^{\pm})\right)_{k\in\mathbb{N}}$. Since $\dim_{\mathrm{H}}mE=\dim_{\mathrm{H}}mE^{\prime}$, Theorem 3.3 gives the following:
Corollary 3.6.
For any $\alpha\in[0,1]$ there exists a distal loom surface having
$\dim_{\mathrm{H}}\overline{Nx_{0}}=1+\alpha.$
3.1.2. Locally Varying Dimensions
In fact, in [SS10] the authors construct compact sets $E^{\prime}\subset[0,1]$ satisfying:
$\alpha_{m}=\dim_{\mathrm{H}}mE^{\prime}\quad,\quad\beta_{m}=\underline{\dim}_{\mathrm{M}}mE^{\prime}\quad,\text{ and}\quad\gamma_{m}=\overline{\dim}_{\mathrm{M}}mE^{\prime},$
where $\overline{\dim}_{\mathrm{M}}$ and $\underline{\dim}_{\mathrm{M}}$ denote the upper and lower Minkowski dimensions, respectively, and where
$0\leq\alpha_{m}\leq\beta_{m}\leq\gamma_{m}\leq 1$
are arbitrary non-decreasing sequences (where $\{\beta_{m}\}$ and $\{\gamma_{m}\}$ are required to satisfy some mild growth constraints). See also [Kör08].
Applying 3.5 and Theorem 3.3 to $E=E^{\prime}+1$, we can produce loom surfaces supporting horocyclic orbit closures having locally varying Hausdorff dimensions and such that the Hausdorff dimension disagrees with the Minkowski dimensions locally, i.e., the sets $\overline{Nx_{0}}\cap B_{r}(a_{t}x_{0})$ have varying different dimensions depending on $t\geq 0$ and r>0r>0.333It is not immediately evident what the local Minkowski dimensions are, as these do not comport well with countable unions. Nevertheless, choosing $\alpha_{m}<\beta_{m}$ implies strict inequalities for $\dim_{\mathrm{H}}<\overline{\dim}_{\mathrm{M}}$.
3.1.3. Discrete Sub-Invariance
Elements of $\Delta_{0}$ correspond to sub-invariance symmetry of $\overline{Nx_{0}}$, that is, $t\in\Delta_{0}$ are exactly those numbers for which
$a_{t}\overline{Nx_{0}}\subset\overline{Nx_{0}},$
see [FLM23, §7]. By constructing $\Sigma_{s}$ with $E=\{T_{0}\}$ for some $T_{0}>0$ we obtain a horocyclic orbit closure with a discrete semigroup of sub-invariance. This is in contrast with orbit closures in $\mathbb{Z}$-covers of compact surfaces, see [FLM23, Prop. 7.20].
3.2. Proof of Theorem 3.3
We proceed with the proof of the main theorem of this section. A key component is Lemma 1.10, which can rephrased in our setting as follows:
Lemma 3.7.
For $j=0,1$ and $t\geq 0$, the point $a_{t}x_{j}$ is contained in $\overline{Nx_{0}}$ if and only if there exists $y_{m}\to x_{j}$ with $A_{+}y_{m}$ asymptotic to $A_{+}x_{0}$ and $\mathscr{S}_{+}(A_{+}y_{m})\to t$.
Proof of Theorem 3.3.
Let us first consider the claim
$\Delta_{0}=\bigcup_{k\in\mathbb{N}}2kE.$
Let $t\in 2kE$ with $t=\sum_{j=1}^{2k}e_{j}$ where $e_{j}\in E$. Let $\varepsilon>0$ and let $S>0$ be the constant in 1.9 corresponding to $\varepsilon$ and $m=2k$. Let $(s_{k})$ be the sequence of horizontal positions of boundary components in the definition of $\Sigma_{s}$. Let $T>0$ be large enough that $|s_{k}-s_{l}|>S$ holds for all distinct $s_{k},s_{l}>T$. By the definition of $E$ there exists for each $e_{j}$ a crossing $\eta^{\pm}_{k_{j}}$ with $|\mathscr{S}_{+}(\eta_{k_{j}}^{\pm})-e_{j}|<\varepsilon/2k$ and with $s_{k_{j}}>T$. Hence the weaving geodesic $\eta_{W}^{+}$ with weaving pattern $\{k_{1},...k_{2k}\}$ satisfies the conditions of 1.9 implying:
$\left|\mathscr{S}_{+}(\eta_{W}^{+})-\sum_{j=1}^{2k}e_{j}\right|<\left|\mathscr{S}_{+}(\eta_{W}^{\pm})-\sum_{j=1}^{2k}\mathscr{S}_{+}(\eta^{\pm}_{k_{j}})\right|+\varepsilon<2\varepsilon.$
Notice that the weaving geodesic $\eta_{W}^{+}$ is both backward and forward asymptotic to $Ax_{0}$. In particular, as $T$ increases the distance between $x_{0}$ and such $\eta_{W}^{+}$ tends to 0. We may thus choose $T$ so large such that $d(x_{0},y)<\varepsilon$ for some $y\in\eta_{W}^{+}$ satisfying $|\mathscr{S}_{+}(A_{+}y)-\mathscr{S}_{+}(\eta_{W}^{+})|<\varepsilon$. Since $\varepsilon>0$ was arbitrary we conclude from Lemma 3.7 that $a_{t}x_{0}\in\overline{Nx_{0}}$, i.e. that $t\in\Delta_{0}$.
In the other direction, let $t\in\Delta_{0}$ then by Lemma 3.7 there exist a sequence of $y_{m}\to x_{0}$ satisfying $\mathscr{S}_{+}(A_{+}y_{m})\to t$. Denote $\rho=t+1$ and let $S>0$ be the constant from 1.6. For all $0<\varepsilon<1$ and all large enough $m$ we have $\mathscr{S}_{+}(A_{+}y_{m})<\rho$ and $d(a_{S+1}y_{m},a_{S+1}x_{0})<\varepsilon$ implying in particular that $\tau(a_{S+1}y_{m})>S$. Hence by 1.6 we conclude that $A_{[S+1,\infty)}y_{m}$ and hence $A_{+}y_{m}$ are weaving geodesic rays. Since $\Sigma_{s}$ is distal we conclude that $A_{+}y_{m}$ has a finite weaving pattern. Moreover, after a possible arbitrarily small perturbation (along a short expanding horocycle) we may assume that $y_{m}$ is backward asymptotic to $Ax_{0}$. Therefore $Ay_{m}=\eta_{W}^{+}$ for some finite even weaving pattern $W$. By identical considerations as before we can approximate $t\approx\mathscr{S}_{+}(A_{+}y_{m})\approx\mathscr{S}_{+}(\eta_{W}^{+})$ by an even sum of slacks of crossings, as claimed.
Similar considerations yield the description of $\Delta_{1}$ in the statement of the theorem. ∎
Appendix A Construction of Invariant Measure
In this section we prove the existence of a locally finite conservative $N$-invariant measure supported on the $N$-minimal set constructed in Theorem 2.2. We first prove a more general proposition concerning continuous flows having a “nice” section, and then move on to apply the proposition to our setting.
Let $X$ be a locally compact second countable Hausdorff space together with a continuous $\mathbb{R}$-action (flow) $\varphi_{t}$. Given a subset $I$ of $\mathbb{R}$ and $F\subseteq X$ we denote $\varphi_{I}F:=\{\varphi_{t}x:t\in I,x\in F\}$. A subset $Y\subseteq X$ is called $\varphi$-minimal if $Y=\overline{\varphi_{\mathbb{R}}y}$ for all $y\in Y$. A (partial) section $\Psi\subseteq X$ for the $\varphi$-flow is a subset satisfying that $\left\{t\in\mathbb{R}:\varphi_{t}(x)\in\Psi\right\}$ is either discrete in $\mathbb{R}$ or empty, for all $x\in X$.
Proposition A.1.
Let $Y\subseteq X$ be a $\varphi$-minimal set consisting of more than one $\varphi$-orbit. Assume there exists a precompact section $\Psi\subseteq X$ satisfying the following:
  1. (1)
    for all $R>0$ the set $\varphi_{(-R,R)}\Psi$ is open in $X$; and
  2. (2)
    $Y\cap\overline{\Psi}=Y\cap\Psi\neq\emptyset$.
Then $Y$ supports a locally finite $\varphi$-invariant conservative measure.
Whenever the set $Y$ is compact then the classical Krylov–Bogolyubov theorem ensures the existence of an invariant probability measure. The proposition above thus deals with the case where $Y$ is non-compact and where the time orbits spend in any compact set may be of zero density.
Proof.
As explained above, we may assume that $Y$ is non-compact. Fix some $y_{0}\in Y$ and denote the open set $B_{R}=\varphi_{(-R,R)}\Psi$ for any $R>0$. We begin by observing that the orbit $\varphi_{\mathbb{R}}y_{0}$ spends an infinite amount of time in $B_{R}$ for any $R>0$. Indeed, since $B_{R}$ is open and intersects $Y$ non-trivially we know by the minimality of $Y$ and the fact that $Y$ is not one orbit, that $\varphi_{\mathbb{R}}y_{0}$ returns in an unbounded set of times to $B_{R}$. Each passage of $\varphi_{\mathbb{R}}y_{0}$ in $B_{R}$ is at least of length $2R$, implying
$\left|\{t\in\mathbb{R}:\varphi_{t}y_{0}\in B_{R}\}\right|=\infty.$
Given $R>0$, consider the following “statistical” probability measures:
$\nu^{R}_{T}=\frac{\int_{-T}^{T}\mathbbm{1}_{B_{R}}(\varphi_{t}y_{0})\delta_{\varphi_{t}y_{0}}dt}{\int_{-T}^{T}\mathbbm{1}_{B_{R}}(\varphi_{t}y_{0})dt}$
where, $\delta_{x}$ denotes Dirac measure at $x\in X$. In other words, $\nu^{R}_{T}$ measures normalized arc-length along $\varphi_{(-T,T)}y_{0}\cap B_{R}$.
Fix $R>0$. We claim the following:
Claim 1: The family of measures $\nu^{R}_{T}$ is asymptotically tight in $B_{R}$, that is, for every $\varepsilon>0$ there exists a compact subset $K_{\varepsilon}\subset B_{R}$ satisfying
$\nu^{R}_{T}(K_{\varepsilon})\geq 1-\varepsilon\quad\text{for all large enough }T>0.$
Given $0<\varepsilon$ set $0<\eta<\frac{1}{4}\varepsilon\cdot R$ and
$K_{\varepsilon}=Y\cap\overline{B_{R-\eta}}.$
By construction $K_{\varepsilon}$ is compact. Assumption (2) of the proposition implies
$K_{\varepsilon}=Y\cap\varphi_{[-R+\eta,R-\eta]}\overline{\Psi}=Y\cap\varphi_{[-R+\eta,R-\eta]}\Psi,$
that is, $K_{\varepsilon}$ is contained in $B_{R}$.
[…]
$\left|\left\{t\in I_{j}:\varphi_{t}y_{0}\in B_{R}\smallsetminus B_{R-\eta}\right\}\right|\leq 4\eta.$
We therefore have
$1-\nu^{R}_{T}(K_{\varepsilon})$ $\leq\nu^{R}_{T}(B_{R}\smallsetminus B_{R-\eta})\leq$
$\leq\frac{1}{\sum_{j}|I_{j}|}\sum_{j}\left|\left\{t\in I_{j}:\varphi_{t}y_{0}\in B_{R}\smallsetminus B_{R-\eta}\right\}\right|\leq$
$\leq\frac{4\eta\cdot m}{2R\cdot(m-2)}<\frac{\varepsilon}{2}\cdot\frac{m}{m-2},$
where m2m-2 comes from omitting up to two $I_{j}$’s bounded by $\pm T$. Since $Y$ was assumed to be non-compact and $B_{R}$ is precompact, we know that $\varphi_{\mathbb{R}}y_{0}$ exits $B_{R}$ infinitely many times. Hence as $T\to\infty$ we have $m\to\infty$, implying
$\nu^{R}_{T}(K_{\varepsilon})\geq 1-\varepsilon$
for all large enough $T>0$, proving claim 1.
Asymptotic tightness implies that for any $R>0$ and sequence $T_{n}\to\infty$ there exists a subsequence for which the measures $\nu^{R}_{T_{n}}$ converge to a probability measure supported on $Y\cap B_{R}$.
Claim 2: Any such limiting measure $\nu$ is invariant along flow lines in $B_{R}$. That is, if $E\subseteq B_{R}$ is a Borel set, $s\in\mathbb{R}$, and $\varphi_{s}E\subseteq B_{R}$, then
$\nu(E)=\nu(\varphi_{s}E).$
This follows from a standard amenability argument. It suffices to show that for any $f\in C_{c}(B_{R})$ with $f\circ\varphi_{s}\in C_{c}(B_{R})$ then $\nu(f\circ\varphi)=\nu(f)$. If $\mathrm{supp}(f)\cap Y=\emptyset$ the claim is trivially true. Otherwise, for any $T>0$ we have
$|\nu^{R}_{T}(f)-\nu^{R}_{T}(f\circ\varphi_{s})|\leq\frac{2s\cdot\|f\|_{\infty}}{\int_{-T}^{T}\mathbbm{1}_{B_{R}}(\varphi_{t}y_{0})dt},$
and since the denominator tends to infinity as $T\to\infty$ the claim follows.
We are ready to begin the construction in earnest. Let $T_{n}\to\infty$ be a sequence for which $\nu^{1}_{T_{n}}$ converges to a probability measure $\mu_{1}$ supported on $Y\cap B_{1}$. Let $T_{n_{k}}$ be a subsequence for which $\nu^{2}_{T_{n_{k}}}$ converges to a probability measure $\tilde{\mu}_{2}$ on $Y\cap B_{2}$. Taking a further subsequence we get a measure $\tilde{\mu}_{3}$ on $Y\cap B_{3}$ and so on ad infinitum. By a diagonal argument we obtain a sequence $T^{\prime}_{n}$ under which
$\nu^{1}_{T^{\prime}_{n}}\to\mu_{1}\quad\text{and}\quad\nu^{\ell}_{T^{\prime}_{n}}\to\tilde{\mu}_{\ell}\quad\text{for all }2\leq\ell\in\mathbb{N}.$
Since each $\tilde{\mu}_{\ell}$ is flow-line invariant inside $B_{\ell}$ and since finitely many $\varphi$-translates of $B_{1}$ cover $B_{\ell}$ we conclude that $\tilde{\mu}_{\ell}(B_{1})>0$ for all $\ell$. Define
$\mu_{\ell}=\frac{1}{\tilde{\mu}_{\ell}(B_{1})}\tilde{\mu}_{\ell}$.
Claim 3: For any $1\leq\ell_{1}<\ell_{2}$
$\mu_{\ell_{2}}|_{B_{\ell_{1}}}=\mu_{\ell_{1}}.$
Note that the inclusion of open sets $B_{\ell_{1}}\subseteq B_{\ell_{2}}$ gives a natural inclusion of $C_{c}(B_{\ell_{1}})$ in $C_{c}(B_{\ell_{2}})$. By the definition of the statistical measures we thus have for any $T>0$ a constant $C(T)>0$ satisfying
(A.1) $\nu^{\ell_{2}}_{T}(f)=C(T)\cdot\nu^{\ell_{1}}_{T}(f)$
for all $f\in C_{c}(B_{\ell_{1}})$. In fact
$C(T)=\frac{\int_{-T}^{T}\mathbbm{1}_{B_{\ell_{1}}}(\varphi_{t}y_{0})dt}{\int_{-T}^{T}\mathbbm{1}_{B_{\ell_{2}}}(\varphi_{t}y_{0})dt}.$
Since the measures on both sides of (A.1) converge along $T^{\prime}_{n}\to\infty$ we conclude that $C(T^{\prime}_{n})\to C$, where $C$ is independent of $f$, and
$\tilde{\mu}_{\ell_{2}}(f)=C\cdot\tilde{\mu}_{\ell_{1}}(f)\qquad\text{for all }f\in C_{c}(B_{\ell_{1}}),$
or
$\tilde{\mu}_{\ell_{2}}|_{B_{\ell_{1}}}=C\cdot\tilde{\mu}_{\ell_{1}}.$
But after normalization both $\mu_{\ell_{1}}$ and $\mu_{\ell_{2}}$ give $B_{1}$ (a subset of $B_{\ell_{1}}$) mass 1, implying the claim.
We are ready to define the $\varphi$-invariant measure $\mu$ as follows:
$\mu(E):=\lim_{\ell\to\infty}\mu_{\ell}(E\cap Y)$
for any Borel set $E\subseteq X$. Claim 3 implies that the above is an increasing limit, therefore ensuring this function is well-defined. Claim 3 further implies $\mu$ is indeed $\sigma$-additive and that
(A.2) $\mu|_{B_{\ell}}=\mu_{\ell}\quad\text{for all }\ell.$
Crucially, the minimality of $Y$ together with assumptions (1)+(2) imply that $\left\{B_{\ell}\right\}_{\ell}$ is an open cover of $Y$. Hence, given any compact set $K\subset X$ there exists some $\ell$ large enough so that the compact set $K\cap Y$ is contained in $B_{\ell}$. As $\mu(B_{\ell})<\infty$, by construction, we thus conclude that $\mu$ is locally finite (and hence Radon, as $X$ is locally compact second countable and therefore also $\sigma$-compact).
Additionally, given any compact set $K$ and any $s\in\mathbb{R}$, there exists $\ell$ so large such that $(K\cup\varphi_{s}K)\cap Y\subset B_{\ell}$. Hence by claim 2 and (A.2), and by the inner regularity of $\mu$, we conclude that $\mu$ is $\varphi$-invariant.
Conservativity of $\mu$ follows from the fact that any locally finite dissipative measure is necessarily supported on properly embedded $\varphi$-orbits. Our assumption that $Y$ contains more than one orbit excludes this possibility. ∎
Existence of appropriate section
We now focus on the particular setting of this paper. Let $\Sigma_{s}$ be a loom surface satisfying the summability condition and let $Y=\overline{Nx_{0}}$ be $N$-minimal as discussed above. We will deduce the existence of a locally finite invariant measure supported on $Y$ from the existence of a certain section to $Y$, that is, a subset $\Psi\subseteq\mathsf{T}^{1}\Sigma_{s}$ for which $\left\{n\in N:nz\in\Psi\right\}$ is a discrete (possibly empty) subset of $N$ for all $z\in\mathsf{T}^{1}\Sigma_{s}$.
Let $U\leq\operatorname{PSL}_{2}(\mathbb{R})$ denote the upper unipotent subgroup generating the expanding horocycle flow on $\mathsf{T}^{1}\Sigma_{s}$, we accordingly have
$N=\left\{n_{s}=\begin{pmatrix}1&0\\ s&1\end{pmatrix}:s\in\mathbb{R}\right\}\qquad,\qquad U=\left\{u_{r}=\begin{pmatrix}1&r\\ 0&1\end{pmatrix}:r\in\mathbb{R}\right\}.$
Given a subset $J\subseteq\mathbb{R}$, we denote $N_{J}=\{n_{s}:s\in J\}$, $U_{J}=\{u_{r}:r\in J\}$, and $A_{J}=\{a_{t}:t\in J\}$.
Lemma A.2.
There exists a section $\Psi$ for the $N$-flow satisfying:
  1. (1)
    for all $R>0$ the set $N_{(-R,R)}\Psi$ is open in $\mathsf{T}^{1}\Sigma_{s}$; and
  2. (2)
    $Y\cap\overline{\Psi}=Y\cap\Psi\neq\emptyset$.
This lemma together with A.1 concludes the proof of the existence of the desired measure.
Proof.
Let $\delta>0$ denote the injectivity radius at $x_{0}$. As follows from our construction of $Y=\overline{Nx_{0}}$ (see Remark 2.6) we know that $a_{\delta/4}x_{0}$ and $a_{-\delta/4}x_{0}$ are not contained in $Y$, hence there exist small neighborhoods $Q_{\pm}$ around $a_{\pm\delta/4}x_{0}$, respectively, which are disjoint from $Y$. Let $0<\eta<\delta/4$ be sufficiently small so that
$a_{\pm\delta/4}U_{(-\eta,\eta)}x_{0}\subset Q_{\pm}.$
Choose any open interval $(c,d)\subseteq(-\eta/2,\eta/2)$ where both $A_{+}u_{c}x_{0}$ and $A_{+}u_{d}x_{0}$ have infinite slack, e.g. by choosing these rays to have lifts in $\mathsf{T}^{1}\mathbb{H}^{2}$ terminating outside the limit set. We define
$\Psi:=A_{(-\delta/4,\delta/4)}U_{(c,d)}x_{0}.$
Recall that $NAU$ corresponds to the open Bruhat cell in $\operatorname{PSL}_{2}(\mathbb{R})$, which means in particular that the multiplication map $N\times A\times U\to NAU$ is a diffeomorphism (see e.g. [Kna02, Lemma 6.44] and §2.3 in [FLM23] for more details). By our choice of constants and the fact that the parameterizations of $U$, NN and $A$ are of unit speed, we conclude that $N_{(-\delta/4,\delta/4)}\Psi\subset B_{\delta}^{\mathsf{T}^{1}\Sigma_{s}}(x_{0})$ hence showing the multiplication map $N_{(-\delta/4,\delta/4)}\times A_{(-\delta/4,\delta/4)}\times U_{(c,d)}\to N_{(-\delta/4,\delta/4)}\Psi$ is a homeomorphism.
Injectivity of the map above implies, in particular, that $\Psi$ is a section for the $N$-flow. Moreover, the fact that $N_{(-R,R)}\Psi$ is open in $\mathsf{T}^{1}\Sigma_{s}$ for all $R<\delta/4$ implies property (1) of the statement.
Now note that
$\partial\Psi=A_{[-\delta/4,\delta/4]}u_{c}x_{0}\cup A_{[-\delta/4,\delta/4]}u_{d}x_{0}\cup a_{-\delta/4}U_{[c,d]}x_{0}\cup a_{\delta/4}U_{[c,d]}x_{0}.$
By our choice of $c$ and $d$ we know that $Y\cap(Au_{c}x_{0}\cup Au_{d}x_{0})=\emptyset$. In addition, $Y\cap a_{\pm\delta/4}U_{[c,d]}x_{0}\subset Y\cap(Q_{-}\cup Q_{+})=\emptyset$. We have thus concluded property (2), and the proof of the lemma. ∎
A.1. Acknowledgments
OL would like to thank FD for the warm hospitality at Rennes University during January 2024. OL would also like to thank Zemer Kosloff, François Ledrappier, George Peterzil, Ariel Rapaport, and Omri Sarig for various helpful discussions.
JF acknowledges the support of the Institut Henri Poincaré (UAR 839 CNRS-Sorbonne Université) and LabEx CARMIN (ANR-10-LABX-59-01) and DFG – Project-ID 281071066 – TRR 191. OL acknowledges the support of ISF grant 957/25. YM acknowledges the support of NSF grant DMS-2005328.

### v3

@@
 Other examples include surfaces supporting horocyclic orbit closures of arbitrary Hausdorff dimension in $(1,2)$.
-The study of horospherical flows on hyperbolic manifolds dates back to Hedlund in the 1930s [Hed36] and has played an important role in the development of modern homogeneous dynamics.
+The study of horospherical flows on hyperbolic manifolds dates back to Hedlund in the 1930s [14] and has played an important role in the development of modern homogeneous dynamics.
 In the finite-volume (and geometrically finite) cases, both the measure-theoretic and topological properties of the flow have been extensively studied, revealing a remarkable degree of rigidity (see, e.g.
-[Fur73, DS84, Bur90, Rob03, Rat91]).
+[12, 8, 4, 30, 28]).
 In contrast, the behavior of horospherical flows in the general geometrically infinite setting remains much less well understood.
@@
 While the existence of other, more intricate, orbit closures was known for decades, none have been described in detail, leaving much mystery as to their potential regularity and rigidity properties (c.f.
-[DS00, CM10, GL17, Mat16, Bel18b, Led97, Led98]).
-In recent works [FLM23, FLM24], an explicit description of all horocyclic orbit closures was given in the setting of $\mathbb{Z}$-covers of compact hyperbolic surfaces.
+[6, 5, 13, 24, 3, 22, 23]).
+In recent works [10, 11], an explicit description of all horocyclic orbit closures was given in the setting of $\mathbb{Z}$-covers of compact hyperbolic surfaces.
 These orbit closures were shown to be highly irregular; their structural features depend in a delicate way on the geometry of the underlying surface.
@@
 In this paper, we provide the first examples of non-trivial minimal horocyclic subsets as well as new fractional dimensional orbit closures.
-As a consequence of our constructions, we provide the first counterexamples to the horospherical infinite measure rigidity phenomenon, which has been observed in a vast variety of settings, where every horospherically invariant ergodic Radon measure is either quasi-invariant under the geodesic flow (à la Babillot–Ledrappier) or supported on a single closed orbit; see [Bur90, Rob03, Sar04, LS07, Sar10, OP17, Sar19, LL22, Lan21, LLLO23] and Remark 2.8.
+As a consequence of our constructions, we provide the first counterexamples to the horospherical infinite measure rigidity phenomenon, which has been observed in a vast variety of settings, where every horospherically invariant ergodic Radon measure is either quasi-invariant under the geodesic flow (à la Babillot–Ledrappier) or supported on a single closed orbit; see [4, 30, 32, 21, 33, 25, 34, 19, 20, 18] and Remark 2.9.
 Main Results
@@
 Recall that a non-empty $N$-invariant closed set $F\subseteq\Sigma$ is called $N$-minimal if all $N$-orbits in $F$ are dense in $F$.
-A characterization of points with dense horocyclic orbits in $\mathcal{E}$ is given by [Ebe77, Dal00] where it was shown that $\overline{Nx}\neq\mathcal{E}$ if and only if the geodesic ray $A_{+}x$ is quasi-minimizing, that is, $d_{\mathsf{T}^{1}\Sigma}(a_{t}x,x)\geq t-c$ for some $c\geq 0$ and every $t\geq 0$.
+A characterization of points with dense horocyclic orbits in $\mathcal{E}$ is given by [9, 7] where it was shown that $\overline{Nx}\neq\mathcal{E}$ if and only if the geodesic ray $A_{+}x$ is quasi-minimizing, that is, $d_{\mathsf{T}^{1}\Sigma}(a_{t}x,x)\geq t-c$ for some $c\geq 0$ and every $t\geq 0$.
 As a consequence, $\mathcal{E}$ is $N$-minimal if and only if $\Gamma$ is convex co-compact.
 Studying the different possible trajectories of quasi-minimizing rays and their “efficiency” has turned out to be key in the analysis of horocyclic orbit closures.
-Drawing on techniques developed in [FLM23, FLM24] and inspired by examples introduced by Alexandre Bellis in [Bel18a, §1.5.1], we provide a recipe for tailoring geometrically infinite surfaces supporting horocycles with prescribed recurrence properties.
+Drawing on techniques developed in [10, 11] and inspired by examples introduced by Alexandre Bellis in [2, §1.5.1], we provide a recipe for tailoring geometrically infinite surfaces supporting horocycles with prescribed recurrence properties.
 Our main results are the following:
@@
   • Our surfaces are extremely sparse; the injectivity radius along all diverging geodesic rays tends to infinity.
-This implies, in particular, that the tameness conditions imposed in [Sar10] to deduce measure rigidity cannot be removed.
-Equivalent geometric conditions appear in [LL22].
+This implies, in particular, that the tameness conditions imposed in [33] to deduce measure rigidity cannot be removed.
+Equivalent geometric conditions appear in [19].
   • The possible non-regularity of orbit closures we construct is quite extreme, allowing us to construct orbit closures having disagreeing Hausdorff and lower/upper Minkowski dimensions; see §3.1.2.
@@
 Remark to the reader about the proof
-While we rely on techniques developed in [FLM23, FLM24], we will only make use of several elementary insights and lemmas from said papers.
+While we rely on techniques developed in [10, 11], we will only make use of several elementary insights and lemmas from said papers.
 Our proof is fairly self-contained and requires no prior knowledge or understanding of the results in the $\mathbb{Z}$-cover setting.
@@
 $d_{\mathbb{H}^{2}}(\partial D_{h_{k}}(s_{k}),\partial D_{h_{k+1}}(s_{k+1}))\to\infty\quad\text{where}\quad k\to\infty.$
-Note that under the above conditions the function $\text{Inj--rad}:\Sigma_{s}\to(0,\infty)$, assigning the injectivity radius at a point, is a proper map.
+Note that for any sequence $0<h_{k}$ uniformly bounded away from $\frac{\pi}{2}$ there exist $s_{k}\to\infty$, sufficiently large and spread apart, for which $s=(s_{k},h_{k})_{k\in\mathbb{N}}$ defines a loom surface.
+Under the above conditions the function $\text{Inj--rad}:\Sigma_{s}\to(0,\infty)$, assigning the injectivity radius at a point, is a proper map.
 In particular, injectivity radius tends to infinity along any diverging geodesic ray.
@@
 As the composition of $1$-Lipschitz maps, $\tau$ is itself $1$-Lipschitz.
-Let $\lambda=p(Ax_{0})\cup p(Ax_{1})$ and $\mathsf{T}_{+}^{1}\lambda=Ax_{1}\cup Ax_{2}$.
+Let $\lambda=p(Ax_{0})\cup p(Ax_{1})$ and $\mathsf{T}_{+}^{1}\lambda=Ax_{0}\cup Ax_{1}$.
 Observe that $\tau$ is strictly contracting away from $\lambda$ and is isometric along each of its components.
@@
 1.3. Slack
-We consider the notion of $\tau$-slack introduced in [FLM24, §3] which measures how “efficiently” a path progresses towards the end at $\infty$:
+We consider the notion of $\tau$-slack introduced in [11, §3] which measures how “efficiently” a path progresses towards the end at $\infty$:
 Definition 1.2.
@@
 Consider the function $\beta:\mathsf{T}^{1}\Sigma_{s}\to[-\infty,\infty)$ defined by
-$\beta(y)=\tau(y)-\mathscr{S}_{+}(A_{+}y)=\lim_{t\to+\infty}\tau(a_{t}y)-t.$
-It is upper semi-continuous, as a decreasing limit of continuous functions, and also $N$-invariant, see [FLM23, Lemma 6.1]. Therefore, for all $y\in\mathsf{T}^{1}\Sigma_{s}$ we have
+(1.2) $\beta(y)=\tau(y)-\mathscr{S}_{+}(A_{+}y)=\lim_{t\to+\infty}\tau(a_{t}y)-t.$
+It is upper semi-continuous, as a decreasing limit of continuous functions, and also $N$-invariant, see [10, Lemma 6.1]. Therefore, for all $y\in\mathsf{T}^{1}\Sigma_{s}$ we have
 (1.3) $\overline{Ny}\subseteq\beta^{-1}([\beta(y),\infty)).$
@@
 All $y\in\mathsf{T}^{1}\Sigma_{s}$ with $\beta(y)>$-$\infty$ satisfy that for all $\varepsilon>0$ there exists $T>0$ such that $d(a_{t}y,\mathsf{T}^{1}_{+}\lambda)<\varepsilon$ for all $t>T$.
-See the first part of the proof of [FLM23, Thm. 3.4] for more details.
+See the first part of the proof of [10, Thm. 3.4] for more details.
 In the case of loom surfaces, this asymptotic behavior ensures all finite slack geodesic rays eventually follow some weaving pattern.
+Definition 1.5.
+A geodesic ray $A_{+}w$ is said to be weaving with pattern $W=\{k_{1}<k_{2}<$$\subseteq\mathbb{N}$ if $t\mapsto\tau(a_{t}w)$ is monotonically increasing and unbounded, for $t\geq 0$, and
+$W=\left\{k\in\mathbb{N}:p(A_{+}w)\cap\partial D_{h_{k}}(s_{k})\neq\emptyset\right\}.$
+Lemma 1.6 (Weaving Lemma).
+In a loom surface $\Sigma_{s}$, weaving geodesics are uniquely determined by their weaving pattern and initial point.
+In addition, for any $\rho\geq 0$ there exists $S>0$ such that any geodesic ray $A_{+}y$ with slack $\mathscr{S}_{+}(A_{+}y)\leq\rho$ and beginning at $\tau(y)>S$ is weaving.
+Proof.
+First note that since both $J_{s}$ copies are convex and simply connected, any geodesic ray passing from $J_{s}\times\{j\}$ through $\partial D_{h_{k}}(s_{k})$ to $J_{s}\times\{j+1\!\!\mod 2\}$ will either stay indefinitely in that side or have its first return to $J_{s}\times\{j\}$ happen through a different component of $\partial J_{s}$.
+Whenever a weaving geodesic has an infinite weaving pattern, monotonicity of $\tau$-values ensures the indices in the pattern appear in strictly increasing order.
+Uniqueness of geodesic representatives in each homotopy class relative endpoints implies such a weaving geodesic is uniquely determined.
+On the other hand, any geodesic ray entirely contained within one $J_{s}\times\{j\}$ and having unbounded $\tau$-values is necessarily asymptotic to $A_{+}x_{j}$.
+We may thus deduce the claim for weaving geodesics with a finite pattern as well.
+For the second claim, given any $\rho\geq 0$ there exists $k_{0}\in\mathbb{N}$ such
+$\rho<\min_{k\geq k_{0}}d_{\mathbb{H}^{2}}(\partial D_{h_{k}}(s_{k}),\partial D_{h_{k-1}}(s_{k-1})).$
+[…]
+A direct consequence of 1.6 is that any finite slack geodesic ray is eventually weaving.
 Given $k\in\mathbb{N}$ we denote by $\eta_{k}^{+}$ the unique geodesic in $\Sigma_{s}$ which is backward asymptotic to $A_{-}x_{0}$ and forward asymptotic to $A_{+}x_{1}$ crossing once from $J_{s}\times\{0\}$ to $J_{s}\times\{1\}$ through $\partial D_{h_{k}}(s_{k})$.
@@
 We similarly have $\eta_{k}^{-}$, the symmetric geodesic passing from $J_{s}\times\{1\}$ to $J_{s}\times\{0\}$ through $\partial D_{h_{k}}(s_{k})$, see Figure 3.
+Symmetry of the construction implies $\eta_{k}^{\pm}\cap\partial D_{h_{k}}(s_{k})=\{s_{k}+ih_{k}\}$.
 Figure 3.
-Notice that by the symmetry of the construction we have $\eta_{k}^{\pm}\cap\partial D_{h_{k}}(s_{k})=\{s_{k}+ih_{k}\}$.
-Definition 1.5.
-A weaving pattern is a any subset of $\mathbb{N}$ thought of as a strictly increasing sequence of indices $W=\{k_{1}<k_{2}<$...\}.
-A weaving geodesic ray with weaving pattern $W$ is an infinite geodesic ray $A_{+}w$ given by pulling tight the following concatenation:
-$\alpha_{0}\ast\eta^{\sigma_{1}}_{k_{1}}\ast\eta^{\sigma_{2}}_{k_{2}}\ast\eta^{\sigma_{3}}_{k_{3}}$...
-where:
-  1. (1)
-    $\sigma_{k}\in\{+,-\}$ are alternating signs;
-  2. (2)
-    $\alpha_{0}$ is the shortest segment connecting $p(w)$ to $\eta^{\sigma_{1}}_{k_{1}}$;
-  3. (3)
-    the $\ast$ notation is interpreted as concatenation of subsegments beginning and ending at the unique intersection points of the paths;
-  4. (4)
-    the weaving pattern is allowed to be finite in which case we append, at the end, an infinite subray of either $A_{+}x_{0}$ or $A_{+}x_{1}$, as needed.
 Given a path $\alpha$, we denote by $\hat{\alpha}$ the geodesic segment given by pulling $\alpha$ tight, that is, the unique geodesic segment in $\alpha$’s homotopy class relative its endpoints (or endpoints at infinity).
 Given a weaving pattern $W$ we define
-$\eta_{W,+}=\widehat{\eta_{k_{1}}^{+}\ast\eta_{k_{2}}^{-}\ast\eta_{k_{3}}^{+}\ast...}$,
-where the alternating signs begin with a ++. We similarly define $\eta_{W,-}$. We will refer to such $\eta_{W}^{\pm}$ as weaving geodesics with weaving pattern $W$.
+(1.4) $\eta_{W,+}=\widehat{\eta_{k_{1}}^{+}\ast\eta_{k_{2}}^{-}\ast\eta_{k_{3}}^{+}\ast...}$,
+where the $\ast$ notation is interpreted as concatenation of subsegments beginning and ending at the unique intersection points of the paths, see Figure 4.
+We similarly define $\eta_{W,-}$ where the alternating signs of the crossing begin with a ‘-’ sign.
+Whenever the weaving pattern is finite the concatenation ends with a full subray of the final crossing.
 Figure 4. Intersection point of two crossings.
-Lemma 1.6 (Weaving Lemma).
-In a loom surface $\Sigma_{s}$, for any $\rho\geq 0$ there exists $S>0$ such that any geodesic ray $A_{+}y$ with slack $\mathscr{S}_{+}(A_{+}y)\leq\rho$ and beginning at $\tau(y)>S$ is weaving.
-Proof.
-First note that since both $J_{s}\times\{j\}$ are convex and simply connected, any geodesic ray passing from $J_{s}\times\{j\}$ through $\partial D_{h_{k}}(s_{k})$ to $J_{j+1}$ will either stay indefinitely in the $j+1\mod 2$-side or will have its first return to $J_{s}\times\{j\}$ happen through a different component of $\partial J_{s}$.
-Therefore, by the uniqueness of geodesic representatives in each homotopy class, it’s enough to show that the sequence of boundary components of $J_{s}$ which the geodesic crosses has increasing indices.
-Now for any $\rho\geq 0$ there exists $k_{0}\in\mathbb{N}$ such
-$\rho<\min_{k\geq k_{0}}d_{\mathbb{H}^{2}}(\partial D_{h_{k}}(s_{k}),\partial D_{h_{k-1}}(s_{k-1})).$
-[…]
-A direct consequence of 1.6 is that any finite slack geodesic ray is eventually weaving.
+The weaving lemma allows us to draw the following realization lemma:
+Lemma 1.7.
+Every weaving geodesic $A_{+}w$ with pattern $W$ is given by pulling tight $\alpha_{0}\ast\eta_{W,\sigma}$ where $\alpha_{0}$ is any path connecting $p(w)$ to $\eta^{\sigma_{1}}_{k_{1}}$ fully contained within one $J_{s}\times\{j\}$ and $\sigma\in\{\pm\}$ is chosen appropriately.
 1.6. Slack of a Weaving Geodesic
@@
 In fact using the hyperbolic cosine law one can verify the following:
-Fact 1.7.
+Fact 1.8.
 $\mathscr{S}_{+}(\eta_{k}^{\pm})=2\ln\cosh(d_{\mathbb{H}^{2}}(0,ih_{k}))=2\ln\cosh\left(\int_{0}^{h_{k}}\frac{dt}{\cos(t)}\right)$.
-We invoke the following lemma from [FLM24, Lemma 3.5]:
+We invoke the following lemma from [11, Lemma 3.5]:
 Lemma 1.9.
@@
 $\left|\mathscr{S}_{+}(\eta_{W,\pm})-\sum_{j=1}^{m}\mathscr{S}_{+}\left(\eta_{k_{j}}^{\pm}\right)\right|<\varepsilon.$
-A fundamental observation, highlighting the utility of the notion of slack, is the following lemma proven in [FLM24, Lemma 3.3]:
+A fundamental observation, highlighting the utility of the notion of slack, is the following lemma proven in [11, Lemma 3.3]:
 Lemma 1.11.
@@
 A loom surface $\Sigma_{s}$ with $s=(s_{j},h_{j})_{j\in\mathbb{N}}$ is said to satisfy the summability condition if
-$\sum_{j=1}^{\infty}\mathscr{S}_{+}(\eta_{j}^{\pm})=\sum_{j=1}^{\infty}2\ln\cosh(d_{\mathbb{H}^{2}}(0,ih_{k}))<\infty.$
+$\sum_{j=1}^{\infty}\mathscr{S}_{+}(\eta_{j}^{\pm})=\sum_{j=1}^{\infty}2\ln\cosh(d_{\mathbb{H}^{2}}(0,ih_{j}))<\infty.$
 We prove the following:
@@
 $Nx_{0}\subsetneq F\subsetneq\mathcal{E},$
-where $\mathcal{E}$ denotes the non-wandering set for the horocycle flow on $\mathsf{T}^{1}\Sigma_{s}$.
+where $\mathcal{E}$ denotes the non-wandering set for the horocyclic flow on $\mathsf{T}^{1}\Sigma_{s}$.
 Recall that two points $y,w\in\mathsf{T}^{1}\Sigma_{s}$ are called $A$-proximal if
-$\liminf_{t\to+\infty}d(a_{t}y,a_{t}w)=0.$
-We will make use of the following fact, proven in [FLM23, Cor. 8.3]:
+$\liminf_{t\to+\infty}d_{\mathsf{T}^{1}\Sigma_{s}}(a_{t}y,a_{t}w)=0.$
+We will make use of the following fact, proven in [10, Cor. 8.3]:
 Proposition 2.3.
 If $y,w\in\mathsf{T}^{1}\Sigma_{s}$ are $A$-proximal then $\overline{Ny}=\overline{Nw}$.
-Our strategy for proving Theorem 2.2 will be to show that all the points in $\overline{Nx_{0}}$ are $A$-proximal to $x_{0}$. Two key observations under these assumptions:
-Lemma 2.4.
-If $\beta(y)=b>$-$\infty$ then $y$ is $A$-proximal to $a_{b}x_{0}$.
+Our strategy for proving Theorem 2.2 will be to show that all the points in $\overline{Nx_{0}}$ are $A$-proximal to $x_{0}$.
+Recall our notation for the projection $p:T^{1}\Sigma_{s}\to\Sigma_{s}$. The half planes
+$\mathcal{O}=\left\{z\in\mathbb{H}^{2}:\mathrm{Im}z<0\right\}\times\left\{0,1\right\},$
+whose boundaries are $p(A_{+}x_{0})$ and $p(A_{+}x_{1})$ respectively, are isometrically embedded in $\Sigma_{s}$.
+Therefore if a geodesic ray $\ell_{+}=p(A_{+}w)$ in $\Sigma_{s}$ enters $\mathcal{O}$ it can never leave it.
+We observe the following:
+Lemma 2.4 (Bottleneck Lemma).
+There exists a sequence $\delta_{k}\to 0$ for which any $w\in\mathsf{T}^{1}\Sigma_{s}$ with $\sup\tau\left(A_{+}w\right)=+\infty$ and $p(A_{+}w)\subseteq\Sigma_{s}\smallsetminus\mathcal{O}$ satisfies
+$d_{\mathsf{T}^{1}\Sigma_{s}}\left(a_{t_{k}}w,a_{s_{k}}x_{0}\right)\leq\delta_{k}\quad\text{for all large }k,$
+where $t_{k}\to\infty$ is the sequence of times in which $\tau(a_{t_{k}}w)=\tau(a_{s_{k}}x_{0})=s_{k}$.
+More generally, if $A_{+}w$ is any geodesic ray with unbounded positive $\tau$-value then
+(2.1) $d_{\mathsf{T}^{1}\Sigma_{s}}\left(a_{t_{k}}w,a_{s_{k}}x_{0}\right)\to 0,$
+for similarly chosen $t_{k}$’s.
+This lemma describes a sequence of arbitrarily thin “bottlenecks”, depicted in Figure 3, which every geodesic ray with unbounded positive $\tau$-value in $\Sigma_{s}\smallsetminus\mathcal{O}$ is forced to pass through.
 Proof.
 First, note that the summability condition implies in particular that $h_{k}\to 0$ and hence
-(2.1) $\lim_{k\to\infty}d\left(a_{s_{k}}x_{0},a_{s_{k}}x_{1}\right)=0.$
-In particular, for all $k$ large enough we have
-(2.2) $\mathsf{T}^{1}_{+}\lambda\cap\tau^{-1}\left((s_{k}-\varepsilon,s_{k}+\varepsilon)\right)\subset B_{3\varepsilon}(a_{s_{k}}x_{0}).$
-Assume $y$ as in the statement.
-For any $\varepsilon>0$ there exists $T>0$ large enough such that $a_{t}y$ is in the $\varepsilon$-neighborhood of $\mathsf{T}^{1}_{+}\lambda$ for all $t\geq T$ (1.4).
-By the definition of $\beta$ we have $|\tau(a_{t-b}y)-t|\to 0$ as $t\to\infty$.
-Hence for all $k$ large enough we have that $|\tau(a_{s_{k}-b}y)-\tau(a_{s_{k}}x_{0})|<\varepsilon$ and $a_{s_{k}-b}y$ is in the $\varepsilon$-neighborhood of $\mathsf{T}^{1}_{+}\lambda$.
-By (2.2) we conclude
-$\liminf_{k\to\infty}d(a_{s_{k}-b}y,a_{s_{k}-b}a_{b}x_{0})\leq 3\varepsilon,$
-implying the claim. ∎
-The second observation is:
+(2.2) $\lim_{k\to\infty}d_{\mathsf{T}^{1}\Sigma_{s}}\left(a_{s_{k}}x_{0},a_{s_{k}}x_{1}\right)=0.$
+Let $I_{k}=[a_{s_{k}}x_{,}a_{s_{k}}x_{1}]$ be the “vertical” geodesic segment connecting the $a_{s_{k}}x_{i}$’s. Note that $I_{k}=\tau^{-1}(s_{k})\cap\Sigma_{s}\smallsetminus\mathcal{O}$.
+Let $w\in\mathsf{T}^{1}\Sigma_{s}$ with $\ell_{+}=p(A_{+}w)\subseteq\Sigma_{s}\smallsetminus\mathcal{O}$ be as in the statement of the lemma.
+For all large $k$, the balls $B_{k}=B_{3}(a_{s_{k}}x_{0})$ of radius 3 around $a_{s_{k}}x_{0}$ are isometrically embedded in $\Sigma_{s}$ and contain $\mathsf{T}^{1}I_{k}=p^{-1}(I_{k})$.
+The geodesic segments $A_{[-1,1]}a_{s_{k}}x_{i}$ are contained in $B_{k}$ for all large $k$ and their Hausdorff distance tends to 0 in $\mathsf{T}^{1}\Sigma_{s}$.
+By our assumption, $\ell_{+}\cap B_{k}\subseteq B_{k}\smallsetminus\mathcal{O}$ implying that $\ell_{+}\cap I_{k}\neq\emptyset$, see Figure 3.
+One can readily verify that
+$\delta_{k}:=\textrm{diam}\left\{y\in\mathsf{T}^{1}I_{k}:p\left(A_{[-1,1]}y\right)\cap p\left(A_{[-1,1]}a_{s_{k}}x_{0}\cup A_{[-1,1]}a_{s_{k}}x_{1}\right)=\emptyset\right\}$
+tends to 0 as $k\to\infty$. Since $p\left(A_{+}w\right)\cap p\left(Aa_{s_{k}}x_{0}\cup Aa_{s_{k}}x_{1}\right)=\emptyset$ we have
+(2.3) $d_{\mathsf{T}^{1}\Sigma_{s}}\left(a_{t_{k}}w,a_{s_{k}}x_{0}\right)\leq\delta_{k}\qquad\text{for all large }k,$
+where $a_{t_{k}}w\in\mathsf{T}^{1}I_{k}$, as claimed.
+More generally, notice that any geodesic ray $A_{+}w$ with unbounded $\tau$-value is either eventually contained in $\Sigma_{s}\smallsetminus\mathcal{O}$ or is asymptotic to one of the $A_{+}x_{i}$’s.
+In both cases we conclude (2.1).
+∎
+As a direct corollary we deduce:
+Corollary 2.5.
+If $\beta(w)=b>$-$\infty$ then $w$ is $A$-proximal to $a_{b}x_{0}$.
+Proof.
+Assume $\beta(w)=b>$-$\infty$, then in particular $\tau(a_{t}w)\to+\infty$ as $t\to+\infty$ and hence (2.1) holds for $\tau(a_{t_{k}}w)=\tau(a_{s_{k}}x_{0})$.
+By the definition of $\beta(w)$, in (1.2), we have
+$\lim_{k\to\infty}s_{k}-t_{k}=\lim_{k\to\infty}\tau(a_{t_{k}}w)-t_{k}=\beta(w)=b,$
+that is, $d_{\mathsf{T}^{1}\Sigma_{s}}\left(a_{t_{k}}w,a_{t_{k}+b}x_{0}\right)\to 0$, as claimed. ∎
+A second useful observation is the following:
 Lemma 2.6.
-For any $\varepsilon>0$ there exists $k_{0}\in\mathbb{N}$ such that any weaving pattern $W$ satisfying $W\geq k_{0}$ will have $\mathscr{S}_{+}(\eta_{W,\pm})<\varepsilon$.
+For any $\varepsilon>0$ there exists $k_{0}\in\mathbb{N}$ for which all weaving patterns $W$ with $W\geq k_{0}$ satisfy $\mathscr{S}_{+}(\eta_{W,\pm})<\varepsilon$.
 Proof.
@@
 $\sum_{j=k_{0}}^{\infty}\mathscr{S}_{+}(\eta_{j}^{\pm})<\varepsilon.$
-Recall that the past of $\eta_{W,\pm}$ is asymptotic to $\mathsf{T}^{1}_{+}\lambda$ and that this geodesic has arbitrarily long segments which are arbitrarily close to $\mathsf{T}^{1}_{+}\lambda$ (close to intersection points of the crossings).
-This is also true for the “untightened” path $\alpha_{W}=\eta_{k_{1}}^{\pm}\ast\eta_{k_{2}}^{\mp}\ast$....
-Hence we can compare slacks along exhausting subsegments of $\eta_{W,\pm}$ and $\alpha_{W}$ which have $\tau$-value arbitrarily close at their respective endpoints.
-The main difference is that the path along $\eta_{W,\pm}$ is shorter.
-Therefore we have $\mathscr{S}_{+}(\eta_{W,\pm})<\mathscr{S}_{+}(\alpha_{W})$.
-The inequality
-$\mathscr{S}_{+}(\alpha_{W})<\sum_{j=k_{0}}^{\infty}\mathscr{S}_{+}(\eta_{j}^{\pm})$
-follows from the additivity and non-negativity of slack. ∎
+Let $W=\{k_{1},k_{2},$...\} be any weaving pattern with $W\geq k_{0}$. By the additivity and non-negativity of slack we conclude that the “untightened” broken path
+$\alpha_{W}=\eta_{k_{1}}^{\pm}\ast\eta_{k_{2}}^{\mp}\ast$...,
+given by concatenating the appropriate subsegments of each crossing has
+$\mathscr{S}_{+}(\alpha_{W})\leq\sum_{j=k_{0}}^{\infty}\mathscr{S}_{+}(\eta_{j}^{\pm})<\varepsilon.$
+The weaving geodesic $\eta_{W,\pm}=\widehat{\alpha_{W}}$ is given by pulling $\alpha_{W}$ tight.
+By construction, both $\alpha_{W}$ and $\eta_{W,\pm}$ are contained in $\Sigma_{s}\smallsetminus\mathcal{O}$ and have unbounded $\tau$-values, implying by Lemma 2.4 the existence of $p^{+}_{k}\in\eta_{W,\pm}$ and $q^{+}_{k}\in\alpha_{W}$ satisfying
+$\tau(p^{+}_{k})=\tau(q^{+}_{k})=s_{k}\to+\infty\qquad\text{and}\qquad d_{\Sigma_{s}}(p^{+}_{k},q^{+}_{k})\to 0.$
+Moreover, both paths are backwards asymptotic to one of the $A_{-}x_{i}$’s giving sequences $p^{-}_{k}$ and $q^{-}_{k}$ with equal $\tau$-values tending to -$\infty$ and having $d_{\Sigma_{s}}(p^{-}_{k},q^{-}_{k})\to 0$.
+Denote $\eta_{W,\pm}^{k}$ the subsegment of $\eta_{W,\pm}$ connecting $p^{-}_{k}$ to $p^{+}_{k}$, and by $\alpha_{W}^{k}$ the subsegment of $\alpha_{W}$ connecting $q^{-}_{k}$ to $q^{+}_{k}$.
+By construction, the length of $\eta_{W,\pm}^{k}$ is shorter than the length of $p^{-}_{k},q^{-}_{k}]\ast\alpha_{W}^{k}\ast[p^{+}_{k},q^{+}_{k}$] implying
+$\mathscr{S}_{+}(\eta_{W,\pm})$ $=\lim_{k\to\infty}\mathscr{S}_{+}(\eta_{W,\pm}^{k})\leq$
+[…]
+$=\mathscr{S}_{+}(\alpha_{W})+0+0<\varepsilon.$
+We are now set to prove the theorem:
 Proof of Theorem 2.2.
-First, by (2.1) we note that $x_{0}$ and $x_{1}$ are $A$-proximal.
-Therefore by 2.3 we have $x_{1}\in F=\overline{Nx_{0}}$.
-Since clearly $x_{1}\notin Nx_{0}$ (the corresponding rays $A_{+}x_{j}$ are not asymptotic) we deduce that $F\neq Nx_{0}$.
-The strict inclusion $F\subsetneq\mathcal{E}$ follows from $x_{0}$ being quasi-minimizing.
-The main part of the claim is the $N$-minimality. Our strategy will be to show that all accumulation points of $Nx_{0}$ are $A$-proximal to $x_{0}$ therefore by 2.3 have dense $N$-orbits in $F$.
-Let $y\in F$ be an accumulation point of $Nx_{0}$ and let $0<\varepsilon<1$ be arbitrary.
-By Facts 1.3 and 1.4 we know $\mathscr{S}_{+}(A_{+}y)<\infty$ and that for any $\varepsilon>0$ a subray of $A_{+}y$ is contained in the $\varepsilon$-neighborhood of $\mathsf{T}^{1}_{+}\lambda$.
-Let $T_{1}>0$ be such that $\mathscr{S}_{+}(A_{[T_{1},\infty)}y)<\varepsilon$ and $A_{[T_{1},\infty)}y$ is contained in the $\varepsilon$-neighborhood of $\mathsf{T}^{1}_{+}\lambda$.
-[…]
-Now let $nx_{0}$ be sufficiently close to $y$ so as to satisfy $d(a_{T_{2}}nx_{0},a_{T_{2}}y)<\varepsilon$, for $T_{2}$ above, and also $d(nx_{0},y)<1$.
-Set $w=a_{T_{2}}nx_{0}$.
-Recall that the function $\beta$ is $N$-invariant and hence $\beta(nx_{0})=\beta(x_{0})=0$ implying that $\mathscr{S}_{+}(A_{+}nx_{0})=\tau(nx_{0})$.
-Since $\tau$ is 1-Lipschitz we have $\tau(nx_{0})<\tau(y)+1$ and hence
-$\mathscr{S}_{+}(A_{+}w)\leq\mathscr{S}_{+}(A_{+}nx_{0})<\tau(y)+1=\rho.$
-By our choice of $T_{2}$ we conclude that $A_{+}w$ is weaving with pattern $W\geq k_{0}$.
-That is, $A_{+}w$ is asymptotic to some $\eta_{W,\pm}$ and moreover $d(w,\eta_{W,\pm})<2\varepsilon$.
-This implies that
-(2.3) $\mathscr{S}_{+}(A_{+}w)<2\varepsilon+\mathscr{S}_{+}(\eta_{W,\pm})<3\varepsilon.$
-But recall that $d(a_{t}nx_{0},a_{t}x_{0})\to 0$ as $t\to\infty$ and hence $d(a_{t-T_{2}}w,a_{t}x_{0})\to 0$. Joining this fact with (2.3) we deduce that $|\tau(w)-T_{2}|<3\varepsilon$.
-Since $d(w,a_{T_{2}}y)<\varepsilon$ we have $|\tau(a_{T_{2}}y)-T_{2}|<4\varepsilon$. On the other hand, by our choice of $T_{2}>T_{1}$ we had
-$\mathscr{S}_{+}(A_{[T_{2},\infty)}y)<\varepsilon.$
-We thus conclude $|\beta(a_{T_{2}}y)-T_{2}|<5\varepsilon$, or in other words, $|\beta(y)|<5\varepsilon$ which by Lemma 2.4 implies $y$ is $A$-proximal to $a_{b}x_{0}$ for some $|b|<5\varepsilon$ (in fact, necessarily $0\leq b$).
-Since $\varepsilon$ was arbitrary we conclude the claim.
-∎
+First, by (2.2) we know that $x_{0}$ and $x_{1}$ are $A$-proximal, therefore implying $x_{1}\in F=\overline{Nx_{0}}$ by 2.3.
+Since $x_{1}\notin Nx_{0}$, as the corresponding $A_{+}x_{j}$’s are not asymptotic, we deduce that $F\neq Nx_{0}$.
+Recall that horocycles outside of $\mathcal{E}$, those who are based outside the limit set, are properly embedded in the quotient manifold, hence $Nx_{0}\subseteq\mathcal{E}$.
+Strict inclusion $F\subsetneq\mathcal{E}$ follows, for instance, from $x_{0}$ being quasi-minimizing, see [9, 7].
+The main part of the claim is the $N$-minimality.
+As promised, our strategy will be to show that all accumulation points of $Nx_{0}$ are $A$-proximal to $x_{0}$ and therefore, by 2.3, have dense $N$-orbits in $F$.
+By 2.5, it suffices to prove that $\beta(y)=0$ for all $y\in F$, which by definition amounts to showing:
+$\mathscr{S}_{+}(A_{+}y)=\tau(y).$
+Let $y\in F$ and let $0<\varepsilon<1$ be smaller than half the minimal injectivity radius in $\Sigma_{s}$.
+By 1.6 and Lemma 2.6 there exists $k_{0}$ large enough so that any geodesic ray $A_{+}w$ beginning at $\tau(w)\geq s_{k_{0}}$ with $\mathscr{S}_{+}(A_{+}w)\leq\rho=\tau(y)+1$ is weaving with pattern $W\geq k_{0}$ for which
+(2.4) $\mathscr{S}_{+}(\eta_{W,\pm})<\varepsilon.$
+By increasing $k_{0}$ if necessary and using Lemma 2.4 we may ensure $\delta_{k}<\varepsilon$ for all $k\geq k_{0}$.
+Now fix $T\geq 0$ for which $d_{\mathsf{T}^{1}\Sigma_{s}}(a_{T}y,a_{s_{k}}x_{0})<\varepsilon$ with $\tau(a_{T}y)=s_{k}>s_{k_{0}}+1$ for some $k>k_{0}$ as ensured by Lemma 2.4.
+By increasing $T$ if necessary we may further require that $\mathscr{S}_{+}(A_{[T,\infty)}y)<\varepsilon$, as this expression decreases to 0 in $T$.
+Let $nx_{0}\in Nx_{0}$ be so close to $y$ so as to satisfy
+(2.5) $d_{\mathsf{T}^{1}\Sigma_{s}}(nx_{0},y)<\varepsilon\quad\text{and}\quad d_{\mathsf{T}^{1}\Sigma_{s}}(a_{T}nx_{0},a_{T}y)<\varepsilon.$
+Denote $w=a_{T}nx_{0}$. The function $\beta$ being $N$-invariant implies $\beta(nx_{0})=\beta(x_{0})=0$ and hence
+(2.6) $\mathscr{S}_{+}(A_{+}nx_{0})=\tau(nx_{0}).$
+We thus have
+$\mathscr{S}_{+}(A_{+}w)$ $\leq\mathscr{S}_{+}(A_{+}nx_{0})=\tau(nx_{0})\leq$
+$\leq\tau(y)+d_{\mathsf{T}^{1}\Sigma_{s}}(nx_{0},y)\leq\tau(y)+1=\rho,$
+by the 1-Lipschitz property of $\tau$.
+Since $\tau(w)>s_{k}-1>s_{k_{0}}$ we are ensured by Lemma 2.6 that $A_{+}w$ is weaving of pattern $W\geq k_{0}$ satisfying (2.4).
+Let $\eta_{W,\sigma}$ correspond to $A_{+}w$ as in Lemma 1.7.
+By Lemma 2.4 and our choice of $a_{T}y$ we know that
+$d_{\Sigma_{s}}(p(w),\eta_{W,\sigma})\leq d_{\Sigma_{s}}(p(w),p(a_{T}y))+d_{\Sigma_{s}}(p(a_{T}y),\eta_{W,\sigma})<\varepsilon+\delta_{k}<2\varepsilon.$
+Let $\alpha$ denote the shortest path connecting $p(w)$ to $\eta_{W,\sigma}$, then $\mathrm{length}(\alpha)<2\varepsilon$ is smaller than the injectivity radius around $p(w)$ implying that $A_{+}w$ is given by pulling tight $\alpha\ast\eta_{W,\sigma}$.
+We thus conclude
+$\mathscr{S}_{+}(A_{+}w)$ $\leq 2\mathrm{length}(\alpha)+\mathscr{S}_{+}(\eta_{W,\sigma})<4\varepsilon+\varepsilon<5\varepsilon,$
+where we made use of the fact that the slack of a path is always bounded above by twice its length, since $\tau$ is $1$-Lipschitz.
+On the other hand, by (2.5) we have
+$|\mathscr{S}_{+}(A_{[0,T]}y)-\mathscr{S}_{+}(A_{[0,T]}nx_{0})|\leq|\tau(y)-\tau(nx_{0})|+|\tau(a_{T}y)-\tau(a_{T}nx_{0})|<2\varepsilon.$
+We therefore obtain
+$|\mathscr{S}_{+}(A_{+}y)-\mathscr{S}_{+}(A_{+}nx_{0})|$ $\leq 2\varepsilon+\mathscr{S}_{+}(A_{[T,\infty)}y)+\mathscr{S}_{+}(A_{[T,\infty)}nx_{0})=$
+$=2\varepsilon+\varepsilon+\mathscr{S}_{+}(A_{+}w)<8\varepsilon.$
+By (2.6) we thus have
+$|\mathscr{S}_{+}(A_{+}y)-\tau(y)|\leq|\mathscr{S}_{+}(A_{+}y)-\mathscr{S}_{+}(A_{+}nx_{0})|+|\tau(nx_{0})-\tau(y)|<9\varepsilon.$
+Since $\varepsilon$ was arbitrary we conclude $\mathscr{S}_{+}(A_{+}y)=\tau(y)$, as claimed. ∎
 Remark 2.7.
-Note that minimality together with (1.2) implies that $\beta$ is constant (zero) on all of $\overline{Nx_{0}}$.
-This also follows directly from our proof.
+We have shown that $\beta$ is constant (zero) on all of $\overline{Nx_{0}}$.
+This also follows from the minimality of $\overline{Nx_{0}}$ together with (1.3).
 Moreover, the fact that $\beta(a_{t}z)=\beta(z)+t$ for all $t\in\mathbb{R}$ implies that $a_{t}y\notin\overline{Nx_{0}}$ for all $y\in\overline{Nx_{0}}$ and $t\neq 0$.
@@
 This theorem provides the first examples of $N$-ergodic and invariant locally finite measures which are neither quasi-invariant with respect to the geodesic flow nor supported on a single closed horocycle.
-Proof of 2.7.
-The existence of an $N$-invariant locally finite measure supported on $\overline{Nx_{0}}$ follows from [KMR13, Lemma 2.2] and the superamenability of the group $\mathbb{R}$ [Ros74].
-See also [Ami22].
+Proof of 2.8.
+The existence of an $N$-invariant locally finite measure supported on $\overline{Nx_{0}}$ follows from [15, Lemma 2.2] and the superamenability of the group $\mathbb{R}$ [31].
+See also [1].
 For completeness and accessibility, we provide a short, more hands-on, construction of such measures in A.1 in the appendix.
@@
 Let $\mu$ denote a locally-finite ergodic component of said measure.
-By Ratner’s classification of finite $N$-invariant measures [Rat92] (and the fact that $\overline{Nx_{0}}$ is not a homogeneous subspace of $\mathsf{T}^{1}\Sigma_{s}$) we conclude $\mu$ is necessarily infinite.
+By Ratner’s classification of finite $N$-invariant measures [29] (and the fact that $\overline{Nx_{0}}$ is not a homogeneous subspace of $\mathsf{T}^{1}\Sigma_{s}$) we conclude $\mu$ is necessarily infinite.
 Singularity of $\mu$ with respect to the geodesic flow follows from the fact that $a_{t}\overline{Nx_{0}}\cap\overline{Nx_{0}}=\emptyset$ for all $t\neq 0$, see Remark 2.7, implying that the topological support of $\mu$ is not invariant under any $a_{t}$ with $t\neq 0$.
@@
 Remark 2.9.
-We note that the notion of a “trivial” measure in this paper is strictly broader than the one used by Sarig in [Sar10], who referred only to measures corresponding to periodic horocycles or orbits based outside the limit set.
-The measures constructed in [Sar10, Theorem 3] are trivial in the sense that they are supported on single proper horocycle orbits, which was known by the author.
-The measures constructed in [Sar10, Theorem 3] are ergodic components of Lebesgue measure on certain surfaces given by Fuchsian groups of the first kind with critical exponent<1/2<\!1/2, see [Pat79].
-In [Pat77, Equation (18)], Patterson shows that Fuchsian groups $\Gamma$ having critical exponent smaller than one-half satisfy, in the Poincaré disk model,
+We note that the notion of a “trivial” measure in this paper is strictly broader than the one used by Sarig in [33], who referred only to measures corresponding to periodic horocycles or orbits based outside the limit set.
+The measures constructed in [33, Theorem 3] are trivial in the sense that they are supported on single proper horocycle orbits, which was known by the author.
+The measures constructed in [33, Theorem 3] are ergodic components of Lebesgue measure on certain surfaces given by Fuchsian groups of the first kind with critical exponent<1/2<\!1/2, see [27].
+In [26, Equation (18)], Patterson shows that Fuchsian groups $\Gamma$ having critical exponent smaller than one-half satisfy, in the Poincaré disk model,
 $\sum_{\gamma\in\Gamma}|\gamma^{\prime}(\xi)|<\infty\quad\text{for Lebesgue-almost every }\xi\in S^{1}.$
 Recall that in the disk model one has $|\gamma^{\prime}(\xi)|=e^{-\beta_{\xi}(\gamma^{-1}.0,0)}$ where $\beta$ is the Busemann cocycle, see e.g.
-[Sar19, Appendix 1].
+[34, Appendix 1].
 Therefore, we have that for Leb-a.e.
@@
 $\Delta_{j}=\{t\geq 0:a_{t}x_{j}\in\overline{Nx_{0}}\cap A_{+}x_{j}\}.$
-See [FLM24, §2.2] and [FLM23, §7] for another account of such sets (under different notations).
-The main theorem in this section is the following:
+See [11, §2.2] and [10, §7] for another account of such sets (under different notations).
+The main theorem of this section, proven in §3.2, is the following:
 Theorem 3.3.
@@
 implying in particular that both $\Delta_{j}$ are closed.
-The vast flexibility of our construction allows us to exhibit any compact subset of $(0,\infty)$ as $E$, that is, we have:
+The vast flexibility of our construction allows us to exhibit any compact [^2] Our definition of loom surfaces is not the most general imaginable.
+In particular, the assumption that the $h_{k}$ are uniformly bounded away from $\frac{\pi}{2}$ is merely imposed for simplicity.
+One can construct surfaces with an unbounded sequence of $h_{k}$ under which additional conditions on $|s_{k+1}-s_{k}|\gg 0$ would ensure all claims in this paper will hold.
+subset of $(0,\infty)$ as $E$, that is, we have:
 Proposition 3.5.
-For any compact [^2] Our definition of loom surfaces is not the most general imaginable.
-In particular, the assumption that the $h_{k}$ are bounded is merely imposed for simplicity.
-One can construct surfaces with an unbounded sequence of $h_{k}$ under which additional conditions on $|s_{k+1}-s_{k}|\gg 0$ would ensure all the claims of this paper will hold.
-set $E\subset(0,\infty)$ there exists a distal loom surface $\Sigma_{s}$ in which $E=\mathrm{accum}\left(\mathscr{S}_{+}(\eta_{k}^{\pm})\right)_{k\in\mathbb{N}}$.
+For any compact set $E\subset(0,\infty)$ there exists a distal loom surface $\Sigma_{s}$ in which $E=\mathrm{accum}\left(\mathscr{S}_{+}(\eta_{k}^{\pm})\right)_{k\in\mathbb{N}}$.
 Proof.
 Given $E$, choose a finite or countable dense subset $\{e_{1},e_{2},$...\} of $E$.
-By solving for $h$ in 1.7 we can construct $\Sigma_{s}$ where every $e_{j}$ appears infinitely many times as the slack of different crossings.
+By solving for $h$ in 1.8 we can construct $\Sigma_{s}$ where every $e_{j}$ appears infinitely many times as the slack of different crossings.
 The proposition follows.
@@
 3.1.1. Fixed Fractional Hausdorff Dimension
-For any $\alpha\in[0,1]$ there exists a compact set $E^{\prime}$ with $\dim_{\mathrm{H}}mE^{\prime}=\alpha$ for all $m\geq 1$, see [SS10, Kör08].
+For any $\alpha\in[0,1]$ there exists a compact set $E^{\prime}$ with $\dim_{\mathrm{H}}mE^{\prime}=\alpha$ for all $m\geq 1$, see [35, 17].
 Since $E^{\prime}$ may contain 0, and the accumulations of slacks of geodesics $\eta_{k}^{\pm}$ are all strictly positive in a distal loom surface, we take $E=1+E^{\prime}$.
@@
 3.1.2. Locally Varying Dimensions
-In fact, in [SS10] the authors construct compact sets $E^{\prime}\subset[0,1]$ satisfying:
+In fact, in [35] the authors construct compact sets $E^{\prime}\subset[0,1]$ satisfying:
 $\alpha_{m}=\dim_{\mathrm{H}}mE^{\prime}\quad,\quad\beta_{m}=\underline{\dim}_{\mathrm{M}}mE^{\prime}\quad,\text{ and}\quad\gamma_{m}=\overline{\dim}_{\mathrm{M}}mE^{\prime},$
@@
 $0\leq\alpha_{m}\leq\beta_{m}\leq\gamma_{m}\leq 1$
-are arbitrary non-decreasing sequences (where $\{\beta_{m}\}$ and $\{\gamma_{m}\}$ are required to satisfy some mild growth constraints). See also [Kör08].
+are arbitrary non-decreasing sequences (where $\{\beta_{m}\}$ and $\{\gamma_{m}\}$ are required to satisfy some mild growth constraints). See also [17].
 Applying 3.5 and Theorem 3.3 to $E=E^{\prime}+1$, we can produce loom surfaces supporting horocyclic orbit closures having locally varying Hausdorff dimensions and such that the Hausdorff dimension disagrees with the Minkowski dimensions locally, i.e., the sets $\overline{Nx_{0}}\cap B_{r}(a_{t}x_{0})$ have varying different dimensions depending on $t\geq 0$ and r>0r>0.333It is not immediately evident what the local Minkowski dimensions are, as these do not comport well with countable unions.
@@
 $a_{t}\overline{Nx_{0}}\subset\overline{Nx_{0}},$
-see [FLM23, §7].
+see [10, §7].
 By constructing $\Sigma_{s}$ with $E=\{T_{0}\}$ for some $T_{0}>0$ we obtain a horocyclic orbit closure with a discrete semigroup of sub-invariance.
-This is in contrast with orbit closures in $\mathbb{Z}$-covers of compact surfaces, see [FLM23, Prop.
+This is in contrast with orbit closures in $\mathbb{Z}$-covers of compact surfaces, see [10, Prop.
 7.20].
@@
 Let $t\in 2kE$ with $t=\sum_{j=1}^{2k}e_{j}$ where $e_{j}\in E$.
-Let $\varepsilon>0$ and let $S>0$ be the constant in 1.9 corresponding to $\varepsilon$ and $m=2k$.
+Let $\varepsilon>0$ and let $S>0$ be the constant in 1.10 corresponding to $\varepsilon$ and $m=2k$.
 Let $(s_{k})$ be the sequence of horizontal positions of boundary components in the definition of $\Sigma_{s}$.
@@
 By the definition of $E$ there exists for each $e_{j}$ a crossing $\eta^{\pm}_{k_{j}}$ with $|\mathscr{S}_{+}(\eta_{k_{j}}^{\pm})-e_{j}|<\varepsilon/2k$ and with $s_{k_{j}}>T$.
-Hence the weaving geodesic $\eta_{W}^{+}$ with weaving pattern $\{k_{1},...k_{2k}\}$ satisfies the conditions of 1.9 implying:
+Hence the weaving geodesic $\eta_{W}^{+}$ with weaving pattern $\{k_{1},...k_{2k}\}$ satisfies the conditions of 1.10 implying:
 $\left|\mathscr{S}_{+}(\eta_{W}^{+})-\sum_{j=1}^{2k}e_{j}\right|<\left|\mathscr{S}_{+}(\eta_{W}^{\pm})-\sum_{j=1}^{2k}\mathscr{S}_{+}(\eta^{\pm}_{k_{j}})\right|+\varepsilon<2\varepsilon.$
@@
 Existence of appropriate section
-We now focus on the particular setting of this paper.
-Let $\Sigma_{s}$ be a loom surface satisfying the summability condition and let $Y=\overline{Nx_{0}}$ be $N$-minimal as discussed above.
-We will deduce the existence of a locally finite invariant measure supported on $Y$ from the existence of a certain section to $Y$, that is, a subset $\Psi\subseteq\mathsf{T}^{1}\Sigma_{s}$ for which $\left\{n\in N:nz\in\Psi\right\}$ is a discrete (possibly empty) subset of $N$ for all $z\in\mathsf{T}^{1}\Sigma_{s}$.
+We now focus on the particular setting of this paper. Let $\Sigma_{s}$ be a loom surface satisfying the summability condition and let $Y=\overline{Nx_{0}}$ be $N$-minimal as discussed above.
 Let $U\leq\operatorname{PSL}_{2}(\mathbb{R})$ denote the upper unipotent subgroup generating the expanding horocycle flow on $\mathsf{T}^{1}\Sigma_{s}$, we accordingly have
@@
 Lemma A.2.
-There exists a section $\Psi$ for the $N$-flow satisfying:
+There exists a section $\Psi\subseteq\mathsf{T}^{1}\Sigma_{s}$ for the $N$-flow satisfying:
   1. (1)
@@
     $Y\cap\overline{\Psi}=Y\cap\Psi\neq\emptyset$.
-This lemma together with A.1 concludes the proof of the existence of the desired measure.
+This lemma together with A.1 concludes the proof of the existence of an $N$-invariant locally finite conservative measure, as desired.
 Proof.
@@
 Recall that $NAU$ corresponds to the open Bruhat cell in $\operatorname{PSL}_{2}(\mathbb{R})$, which means in particular that the multiplication map $N\times A\times U\to NAU$ is a diffeomorphism (see e.g.
-[Kna02, Lemma 6.44] and §2.3 in [FLM23] for more details).
+[16, Lemma 6.44] and §2.3 in [10] for more details).
 By our choice of constants and the fact that the parameterizations of $U$, NN and $A$ are of unit speed, we conclude that $N_{(-\delta/4,\delta/4)}\Psi\subset B_{\delta}^{\mathsf{T}^{1}\Sigma_{s}}(x_{0})$ hence showing the multiplication map $N_{(-\delta/4,\delta/4)}\times A_{(-\delta/4,\delta/4)}\times U_{(c,d)}\to N_{(-\delta/4,\delta/4)}\Psi$ is a homeomorphism.